### v1

# TELL-TALE: Task Efficient LLMs with Task Aware Layer Elimination

Abstract
In this paper we introduce Tale, Task-Aware Layer Elimination, an inference-time algorithm that prunes entire transformer layers in an LLM by directly optimizing task-specific validation performance. We evaluate Tale on 9 tasks and 5 models, including LLaMA 3.1 8B, Qwen 2.5 7B, Qwen 2.5 0.5B, Mistral 7B, and Lucie 7B, under both zero-shot and few-shot settings. Unlike prior approaches, Tale requires no retraining and consistently improves accuracy while reducing computational cost across all benchmarks. Furthermore, applying Tale during finetuning leads to additional performance gains. Finally, Tale provides flexible user control over trade-offs between accuracy and efficiency. Mutual information analysis shows that certain layers act as bottlenecks, degrading task-relevant representations. Tale’s selective layer removal remedies this problem, producing smaller, faster, and more accurate models that are also faster to fine-tune while offering new insights into transformer interpretability.
1 Introduction
While Large Language Models (LLMs) have achieved remarkable success, they come with large computational overheads. This may prevent users in resource limited organizations or with high-throughput applications from using more capable models. This limitation has spawned an important research area on model compression. We focus here on the branch known as model pruning. Although it often reduces computational overhead, pruning typically requires complex implementation procedures, extensive retraining or fine-tuning, and may result in unpredictable performance degradation (Section 2 details related work).
In contrast, having observed that not all layers contribute to a particular task, we propose a lightweight, greedy, iterative layer pruning algorithm, Task Aware Layer Elimination (Tale). Tale operates at inference time, is hardware agnostic, directly optimizes for task-specific accuracy at each pruning step and consistently offers improved results over the original model. This improvement persists in interactions with fine tuning on our tasks. As illustrated in Figure 1 and detailed in Section 3, Tale systematically evaluates all possible single-layer removals at each iteration, selecting the layer whose elimination results in the highest validation accuracy. This process continues iteratively until performance improvements fall below a predefined threshold, ensuring that only layers with minimal or negative impact on task performance are removed.
Tale improves task specific accuracy and provides moderate computational reductions with minimal implementation complexity. This can be extremely useful for agentic systems with LLMs for dedicated tasks; TALE will make those LLMs perform better and more quickly. Our layer-wise pruning leverages the modular nature of transformer architectures, where each layer performs a complete transformation of the input representation through attention and feedforward mechanisms. This architectural property enables the removal of entire layers without requiring modifications to the remaining network structure.
We evaluate Tale with five LLMs, LLaMA 3.1 8B, Qwen 2.5 7B, Qwen 2.5 0.5B, Mistral 7B and Lucie 7B, on 9 diverse benchmark datasets (Sections 4 and 5) both in zero-shot and few-shot settings. We also investigated interactions between pruning with Tale finetuning. We find consistent improvements in both accuracy and computational efficiency in all cases. This suggests pre-trained LLMs contain redundant or even detrimental layers for a given downstream task. We analyze layer flow using the notion of mutual information (MI) to support this hypothesis (Section 6). Additionally, our experiments show Tale’s potential as a tool for understanding layer function in and across models (Section 7), thus aiding model interpretability.
2 Related Work
Zhu et al. (2024) distinguishes four primary approaches to model compression: model pruning, quantization, low-rank approximation, and knowledge distillation. Our work focuses on pruning, which can be categorized into unstructured, structured, and semi-structured methods. Unstructured pruning removes individual parameters, resulting in irregular, sparse structures Han et al. (2015b); Chen et al. (2015); Srinivas & Babu (2015); structured pruning eliminates entire components such as neurons, attention heads, or layers while maintaining the overall network structure He et al. (2017); Voita et al. (2019); Lagunas et al. (2021); Men et al. (2024). Semi-structured pruning combines fine-grained control with structural regularity, and has been explored in recent work Li et al. (2023); Frantar & Alistarh (2023b); Sun et al. (2024). Early pruning methods leveraged second-order information for structured pruning LeCun et al. (1989); Hassibi et al. (1993), but the field has since shifted toward computationally simpler, magnitude-based approaches that prune parameters by importance scores Han et al. (2015a); See et al. (2016); Narang et al. (2017)
Model pruning has benefited from information-theory. Building on the information bottleneck view of deep networks, Tishby et al. (2000); Tishby & Zaslavsky (2015) conceptualize layers as filters that preserve task-relevant information while discarding irrelevant input. This has led to model pruning proposals based on mutual information (MI). Fan et al. (2021) propose a layer-wise strategy that leverages MI estimates to reduce hidden dimensionality in a top-down manner. MI-guided approaches demonstrate that information-theoretic criteria can balance representation quality with model compression (Ganesh et al., 2020; Westphal et al., 2024). A central challenge, however, is the difficulty of estimating MI. Neural estimators such as MINE Ishmael Belghazi et al. (2018) provide principled solutions for high-dimensional settings, but in practice, probing classifiers Belinkov (2022) remain the dominant tool due to their efficiency and interpretability.
With regards to large transformers, Zhang & Papyan (2025) proposes a pruning strategy based on matrix approximations. Xia et al. (2023) demonstrates that pruning 7B models into 1.3B and 2.7B variants via structured layer and hidden-dimension pruning, combined with dynamic batch loading for fine-tuning, can yield submodels outperforming comparably sized models trained from scratch, though not matching the original model. Complementary work Frantar & Alistarh (2023a); Zhang et al. shows that contiguous blocks, especially attention layers, can often be pruned with limited performance loss, although critical “cornerstone” layers Zhang et al. remain essential for reasoning. Kim et al. (2024) investigates lock-level pruning based on weight importance, but these methods typically require fine-tuning techniques to recover baseline accuracy.
Most pruning methods compress models while attempting to match the original performance but often suffer accuracy degradation, which requires retraining or fine-tuning Xia et al. (2024). Improvements are generally measured relative to small models rather than original unpruned baselines. To the best of our knowledge, Tale is the only pruning method reported to improve over the original model’s performance without requiring any additional training or fine-tuning.
Figure 1: Illustration of Tale layer elimination. Candidate layers (yellow) are tested for removal, and the best-performing ones above the threshold are permanently dropped (red) until no further improvement is possible.
3 Tale , our Greedy-Selection Algorithm
3.1 Basics and Intuitions
A transformer maps a sequence of input vectors $(x_{1},\cdots,x_{n})$ to a corresponding sequence of output vectors through a stack of L layers. Each layer $\ell$ transforms the hidden representations $X^{(\ell)}=(x_{1}^{(\ell)},\ldots,x_{n}^{(\ell)})$ into $X^{(\ell+1)}$ through attention and feedforward blocks, connected by residual pathways. Removing layer $\ell$ from this pipeline simply redirects the flow such that $X^{(\ell-1)}\rightarrow X^{(\ell+1)}$, a property that makes the architecture naturally amenable to layer-wise pruning.
Our initial intuition for Tale came from examining the behavior of partial forward passes. Let $h^{(k)}$ denote the hidden representation after $k$ layers. Instead of always decoding from the final representation $h^{(\text{L})}$, we projected intermediate representations $h^{(k)}$ for $k<\text{L}$ directly into the vocabulary space using the output projection $W_{\text{out}}$, i.e.,
$\hat{y}^{(k)}=\mathrm{softmax}(W_{\text{out}}h^{(k)}).$
We then compared the performance of $\hat{y}^{(k)}$ across different values of $k$. Surprisingly, we observed that for many tasks, intermediate layers ($k<\text{L}$) achieved higher accuracy than the final layer L (Figure 5). This indicated that additional depth does not always translate into better task-specific performance: some layers contribute marginally, while others introduce representational noise.
This experiment motivated our central hypothesis: not all layers in an LLM are equally useful, and selectively removing redundant layers can preserve—or even improve—downstream accuracy. Tale (Task-Aware Layer Elimination) formalizes this intuition into a principled, iterative pruning strategy.
3.2 Tale
Tale is a greedy iterative layer pruning algorithm for pre-trained open-weights LLM compression that systematically removes layers while preserving or even improving model performance (Algorithm 6). Starting with a full pre-trained model, Tale evaluates all possible single-layer removals at each iteration, computing the validation accuracy for each candidate pruned architecture. The layer whose removal results in the highest accuracy is permanently eliminated from the model, and this compressed architecture becomes the baseline for the next iteration. This process continues iteratively until the performance improvement falls below a predefined threshold, at which point the algorithm terminates and returns the most compressed model that maintains performance above the specified threshold. Our approach directly optimizes for task-specific accuracy at each pruning step, ensuring that only layers with minimal impact on the target objective are removed. This exhaustive evaluation strategy, while computationally intensive during the pruning phase, provides strong empirical guarantees about the optimality of each pruning decision within the greedy framework.
Algorithm 1 Tale : Greedy Iterative Layer Pruning
1:Pre-trained model $\mathcal{M}$ with $L$ layers; validation set $\mathcal{D}_{val}$; performance threshold $\epsilon$
2:Compressed model $\mathcal{M}^{*}$
3:Initialize $\mathcal{M}^{*}\leftarrow\mathcal{M}$
4:repeat
5:  for each layer $\ell\in\{1,\dots,L\}$ of $\mathcal{M}^{*}$ do
6:   Construct candidate model $\mathcal{M}_{-\ell}$ by removing layer $\ell$
7:   Compute validation accuracy $A_{\ell}=\text{Acc}(\mathcal{M}_{-\ell},\mathcal{D}_{val})$
8:  end for
9:  Select $\ell^{*}=\arg\max_{\ell}A_{\ell}$
10:  if $A_{\ell^{*}}\geq\text{Acc}(\mathcal{M}^{*},\mathcal{D}_{val})-\epsilon$ then
11:   Update $\mathcal{M}^{*}\leftarrow\mathcal{M}_{-\ell^{*}}$
12:  else
13:   break
14:  end if
15:until All Accuracies below threshold
16:return $\mathcal{M}^{*}$
4 Benchmarks and Datasets
We evaluate Tale across a diverse suite of nine benchmarks spanning reasoning, language understanding, and commonsense knowledge. For mathematical reasoning, we include GSM8K-Hard, a curated subset of GSM8K Cobbe et al. (2021) with more than five premises per question to increase difficulty, and MATH500 Hendrycks et al. (2021b), a benchmark for symbolic and arithmetic reasoning (for evaluation details see Appendix A). For language understanding, we consider MMLU Hendrycks et al. (2021a) and BoolQ Clark et al. (2019), while Winogrande Sakaguchi et al. (2021), CommonsenseQA Talmor et al. (2019), and BIG-Bench Srivastava et al. (2023) capture commonsense and multi-task generalization. Finally, we include both ARC-Easy and ARC-Challenge Clark et al. (2018), which evaluate scientific and factual reasoning at varying difficulty levels. Together, these nine datasets cover a broad spectrum of downstream challenges and allow us to assess both the generality and task-specific benefits of our pruning strategy.
5 Results
We evaluate Tale across five medium-scale models (LLaMA 3.1 8B, Mistral 7B, Lucie 7B, Qwen 2.5 7B) and one smaller model (Qwen 2.5 0.5B), spanning nine benchmarks that cover commonsense reasoning, reading comprehension, and mathematical problem solving. All experiments are conducted in the zero-shot setting unless otherwise noted. [^1] Code available at https://anonymous.4open.science/r/tale/ Table 3 summarizes model configurations.
(a) ARC-Easy
(b) ARC-Challenge
(c) BoolQ
(d) MMLU
(e) CommonQA
(f) WinoGrande
(g) BIG-Bench
(h) GSM8K-Hard
(i) Math500
Figure 2: Accuracy progression of Tale across 9 benchmark datasets for LLaMA 3.1 8B. Each curve represents the accuracy at successive iterations. The $\star$ denotes the best-performing layer drop configuration, while the $\square$ highlights the Best Speed up with at least Baseline Accuracy (BSBA) configuration.
Dataset LLaMA 3.1 8B (zero-shot) Qwen 2.5 7B (zero-shot)
Baseline Best Model BSBA Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 87.00 90.55(+4.08% $\uparrow$) 5 1.3 87.82 8 1.4 90.04 94.40(+4.84% $\uparrow$) 3 1.2 90.08 7 1.5
ARC-Challenge 75.86 78.62(+3.63% $\uparrow$) 4 1.3 76.90 7 1.4 86.55 92.00(+6.45% $\uparrow$) 2 1.3 86.55 6 1.4
BoolQ 85.00 86.20(+1.40% $\uparrow$) 3 1.1 85.70 7 1.4 81.90 83.90(+2.44% $\uparrow$) 4 1.3 82.70 5 1.2
MMLU 54.87 59.90(+9.17% $\uparrow$) 1 1.1 54.87 9 1.4 68.10 71.00(+4.26% $\uparrow$) 5 1.2 68.13 6 1.3
CommonQA 72.20 75.30(+4.29% $\uparrow$) 3 1.2 73.10 6 1.3 80.30 84.40(+5.11% $\uparrow$) 2 1.1 80.50 6 1.3
Winogrande 53.83 56.67(+5.28% $\uparrow$) 4 1.3 53.83 12 1.5 62.04 67.25(+8.40% $\uparrow$) 3 1.2 62.19 6 1.5
BIG-Bench 75.20 83.60(+11.17% $\uparrow$) 5 1.3 75.20 11 1.6 79.20 81.60(+3.03% $\uparrow$) 6 1.5 81.60 6 1.5
GSM8K-HARD 15.07 37.08(+146.05% $\uparrow$) 1 1.1 17.31 6 1.4 7.8 26.8(+243.58% $\uparrow$) 2 1.1 8.99 5 1.2
Math500 20.50 26.00(+26.83% $\uparrow$) 1 1.1 26.00 3 1.2 18.00 27.00(+50.0% $\uparrow$) 2 1.1 21.00 4 1.2
Dataset Lucie 7B (zero-shot) Mistral 7B (zero-shot)
Baseline Best Model BSBA Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 72.45 76.55(+5.66% $\uparrow$) 6 1.3 72.55 13 1.8 81.02 83.45(+4.23% $\uparrow$) 5 1.3 81.09 9 1.5
ARC-Challenge 48.00 53.79(+12.06% $\uparrow$) 7 1.4 51.38 11 1.8 72.20 74.83(+3.64% $\uparrow$) 6 1.4 72.41 8 1.4
BoolQ 53.70 77.50(+44.32% $\uparrow$) 5 1.3 60.60 19 2.8 80.36 83.20(+3.53% $\uparrow$) 6 1.2 80.60 10 1.5
MMLU 21.36 42.98(+101.2% $\uparrow$) 8 1.4 22.73 22 3.3 52.73 57.81(+9.63% $\uparrow$) 2 1.0 52.91 8 1.4
CommonQA 55.50 69.70(+25.59% $\uparrow$) 3 1.5 57.10 17 2.6 57.32 61.40(+7.12% $\uparrow$) 4 1.2 57.40 7 1.3
Winogrande 54.20 57.80(+6.64% $\uparrow$) 5 1.2 54.30 15 1.9 52.55 58.80(+11.53% $\uparrow$) 10 1.6 53.43 13 1.8
BIG-Bench 69.60 77.20(+9.84% $\uparrow$) 9 1.2 72.00 15 1.6 70.00 76.40(+9.14% $\uparrow$) 9 1.3 72.80 11 1.4
GSM8K-HARD 14.20 17.80(+25.35% $\uparrow$) 1 1.1 17.40 3 1.1 11.24 19.10(+69.92% $\uparrow$) 2 1.1 15.73 4 1.2
Math500 19.00 27.00(+42.11% $\uparrow$) 2 1.1 26.00 3 1.2 8.00 16.00(+100% $\uparrow$) 1 1.0 10.00 4 1.2
Dataset Qwen 2.5 0.5B (zero-shot)
Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 40.00 60.91(+48.49% $\uparrow$) 3 1.1 48.36 5 1.4
ARC-Challenge 35.52 40.34(+13.57% $\uparrow$) 1 1.1 37.24 4 1.5
BoolQ 62.30 67.20(+7.87% $\uparrow$) 5 1.4 66.20 6 1.5
MMLU 31.48 39.97(+26.96% $\uparrow$) 2 1.1 33.90 5 1.4
CommonQA 42.40 49.10(+15.80% $\uparrow$) 2 1.3 44.00 3 1.4
Winogrande 49.86 51.88(+4.51% $\uparrow$) 5 1.3 49.87 17 3.9
BIG-Bench 72.40 73.60(+1.66% $\uparrow$) 2 1.2 73.60 2 1.2
GSM8K-HARD 6.74 11.24(+66.77% $\uparrow$) 1 1.2 8.99 2 1.2
Math500 8.00 12.00(+50% $\uparrow$) 1 1.1 9 2 1.1
Table 1: Performance comparison across language models under 0-shot evaluation. We report accuracy (Perf.), number of dropped layers (#D), and relative inference speedup (Sp.). Percentage gain = $\frac{\text{Acc}_{\text{Best}}-\text{Acc}_{\text{Baseline}}}{\text{Acc}_{\text{Baseline}}}\times 100$. Best accuracy is highlighted in bold; BSBA shows balanced trade-offs.
Iterative pruning trajectories. Figure 2 visualizes the iterative layer-pruning process for LLaMA 3.1 8B. Each curve tracks accuracy as layers are progressively removed. The $\star$ denotes the best-performing pruned model (Best), while the $\square$ highlights the Best Speedup with Baseline Accuracy (BSBA) model—the pruned configuration achieving maximum inference speedup without falling below baseline accuracy.
From these trajectories, three consistent patterns emerge: (i) Tale identifies compressed models that outperform the original across diverse tasks, with $\star$ markers lying strictly above baseline. (ii) Accuracy improvements persist across multiple pruning steps before diminishing returns, showing that substantial redundancy exists even in carefully tuned pretrained models. (iii) Pruning dynamics are task-specific: datasets such as ARC-Easy and MMLU tolerate deeper pruning while continuing to improve, whereas reasoning-heavy tasks like GSM8K-Hard converge earlier, reflecting heterogeneous layer importance across domains.
Best vs. BSBA models. Table 1 compares baseline models against their pruned counterparts under both Best and BSBA configurations. Across all benchmarks, the Best models yield consistent accuracy gains—up to $146\%$ (LLaMA 8B on GSM8K-Hard), $101\%$ (Lucie 7b on MMLU) and $244\%$ (Qwen 7B on GSM8k-Hard)—while also delivering moderate speedups. BSBA models, by construction, trade smaller gains in accuracy for more aggressive speedups, offering practical operating points when inference cost is the dominant concern.
Few-shot setting. We tested Tale under the few-shot regime for Lucie and LLaMA models (Appendix Tables 45). Few-shot prompting improves baselines on reasoning tasks such as GSM8K and Math500, yet Tale-pruned variants still achieve higher accuracy in nearly all settings. This shows that pruning-induced improvements are largely complementary to gains from in-context learning.
Takeaways. Tale consistently uncovers Pareto-optimal models along the accuracy–efficiency frontier. By balancing task fidelity with computational savings, it enables both accuracy-focused and efficiency-focused deployment. The observed diversity in pruning profiles across datasets underscores the importance of adaptive pruning, rather than one-size-fits-all heuristics, for effective model compression. A tunable selection metric for choosing among candidate trade-offs is in Appendix E.
5.1 Tale and Fine-tuning: How Does Pruning Interact with Training?
A natural question is whether pruning layers before or after fine-tuning harms the model’s ability to learn. Intuitively, one might expect that removing layers reduces representational capacity and thus limits downstream fine-tuning performance compared to baseline instruct-tuned models. Surprisingly, our experiments show the opposite: Tale not only preserves fine-tuning efficacy but in several cases improves both accuracy and efficiency.
We systematically explored four settings: (i) fine-tuning the base model (FT), (ii) applying Tale after fine-tuning (FT $\rightarrow$ Tale ), (iii) pruning first and then fine-tuning (Tale $\rightarrow$ FT), and (iv) pruning first then finetuning and then again pruning (BASE $\rightarrow$ Tale $\rightarrow$ FT $\rightarrow$ Tale ). Across reasoning and knowledge benchmarks, we consistently observed moderate but significant gains, especially on Winogrande (Table 2).
For example, pruning LLaMA-3.1 8B before fine-tuning reduced fine-tuning time by 2–2.5 GPU hours on an A100 (an 18.5% reduction) while simultaneously improving Winogrande performance by $2.4\%$. Iteratively applying pruning and fine-tuning allowed us to prune up to 8 layers achieving still higher accuracy (87.37%) than the full fine-tuned model (85.00%). Similarly, pruning the fully fine-tuned model yielded a 7-layer reduction while maintaining strong accuracy (86.66%).
On knowledge-heavy tasks such as MMLU and CommonsenseQA, the gains were smaller but consistent: pruned fine-tuned models retained or slightly exceeded the accuracy of their full fine-tuned counterparts, while requiring fewer parameters to optimize. This suggests that pruning can act as a regularizer, simplifying the optimization landscape by removing redundant layers.
Model Dataset Baseline Pruned Only FT Only Prune $\rightarrow$ FT FT $\rightarrow$ Prune (Prune $\rightarrow$ FT) $\rightarrow$ Prune
Perf. #D Perf. #D Perf. #D Perf. #D Perf. #D Perf. #D
Llama 3.1 8B Winogrande 53.83 0 56.67 4 85.00 0 87.06 4 86.74 7 87.37 8
MMLU 54.87 0 59.90 1 63.62 0 63.49 1 64.21 2 64.01 2
CommonQA 72.20 0 75.30 3 81.88 0 81.80 3 83.40 3 82.90 6
GSM8K 15.07 0 37.08 3 42.70 0 53.96 1 50.86 2 54.02 2
Qwen 0.5B Winogrande 49.86 0 51.88 5 50.43 0 50.43 5 50.49 2 52.49 9
MMLU 31.48 0 39.98 2 44.87 0 43.76 2 45.53 2 45.58 3
Table 2: Comparison of Llama 3.1 8B and Qwen 0.5B across Winogrande, MMLU, and CommonQA under different pruning and fine-tuning regimes. Columns denote: (i) Baseline = original model, (ii) Pruned Only = Tale without fine-tuning, (iii) FT Only = fine-tuned without pruning, (iv) Prune $\rightarrow$ FT = prune then fine-tune, (v) FT $\rightarrow$ Prune = fine-tune then prune, (vi) (Prune $\rightarrow$ FT) $\rightarrow$ Prune = best fine-tuned-pruned model further pruned. Perf. = performance score, #D = number of deleted layers.
Overall, these results highlight an unexpected but consistent trend: pruning with Tale does not hinder fine-tuning but instead synergizes with it. Pruned models fine-tune faster, require fewer parameters to adapt, and are close to or better in performance than their full counterparts. Pruning finetuned models almost always improves performance. This shows that Tale pruning can be effectively interleaved with fine-tuning to create models that are both more accurate and computationally efficient.
6 Information theory and why pruned models perform better.
Our experiments show that selectively dropping layers improves accuracy for a targeted task. This seems counterintuitive: why could removing parts of a carefully trained model improve it?
Alemi et al. (2016); Tishby & Zaslavsky (2015) have used information theory (Shannon, 1948) to analyze how neural networks learn and represent data. We can use also this tool to explain the effects of our pruning algorithm. Fano & Hawkins (1961) define $\text{I}(\text{X};\text{Y})$, the mutual information between two random variables $X$ and $Y$, with the equation:
$\text{I}(\text{X};\text{Y})=\text{H}(\text{Y})-\text{H}(\text{Y}\mid\text{X})=\text{H}(\text{X})-\text{H}(\text{X}\mid\text{Y})=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\log\frac{p(x,y)}{p(x)\,p(y)}$ (1)
where $p(x,y)$ is the joint distribution of X and Y, and $p(x),p(y)$ are their marginals and where $\text{H}(\text{X})=-\sum_{x}p(x)\log p(x)$ is the Shannon (1948) entropy. $\text{I}(\text{X};\text{Y})$ measures how much knowing X reduces uncertainty about Y Tishby & Zaslavsky (2015); Shwartz-Ziv & Tishby (2017).
To apply these notions in our case, we estimate $\text{I}(\text{X}^{(l)},\text{Y})$ at each layer using trained probes as mutual information approximators. While this approach yields $\text{I}_{\text{probe}}(\text{X}^{(l)},\text{Y})$, an approximation of the true mutual information, probe-based estimates capture the linearly accessible task-relevant information at each layer, which is meaningful since many downstream tasks employ linear classification heads Belinkov (2022).
This estimate together with these three observations yields an information-theoretic explanation for why architectural pruning can improve model performance. (i) Certain layers in large pre-trained transformers decrease mutual information between the layer’s representation and the target task (Figure 3 and Table 7). (ii) From Tale we select the layer $\ell$ removed at the first iteration, and compare $\mathrm{I}\!\left(X^{(\ell+1)},Y\right)$ before deletion of layer $\ell$ to its value after deletion. Deleting this layer always increases the mutual information at the subsequent layer on the tasks, effectively preventing the previous layer, which acts as noise, from obstructing the flow of information through the network. Although our findings depend on the approximative nature of probe-based MI estimation, they provide evidence that certain layers in over-parameterized transformers act as information bottlenecks, which degrade rather than refine task-relevant representations. By removing layers that decrease mutual information with the target task, we enable representations with higher task-relevant information to flow directly to subsequent layers using the residual connection. This improves the information flow through the remaining architecture, results in representations that are more predictive of the target task, and yields improved accuracy at lower computational cost.
(a) ARC-Easy (Qwen 0.5B)
(b) BoolQ (Lucie 7B)
(c) BigBench (Llama 8B)
Figure 3: Evolution of mutual information (MI) across transformer layers for different benchmark datasets and different models. Each subplot shows how information is processed and transformed as it flows through the network layers, demonstrating distinct patterns of information propagation for (a) ARC-Easy on Qwen 0.5B, (b) BoolQ on Lucie 7B, and (c) BigBench on LLama 8B.
7 Discussion
We summarize five key observations below from our experiments.
1. Deleting later layers frequently improves performance on various tasks.
This challenges prior claims that early layers are largely redundant and more amenable to removal, instead highlighting their essential role in preserving core task-relevant representations. Even deleting many late layers does not reduce accuracy below baseline, whereas removing even a single early layer is often catastrophic (see Figure 4). All models exhibit similar behavior.
Figure 4: Nine benchmark tasks indicating performance after one layer is dropped from different positions in Llama3-8B.
On the other hand, early layers appear crucial for task-solving, even though probing outputs at those layers does not yield interpretable responses. These results are potentially important for model interpretability. Plotting the performance degradation from ablating individual or consecutive layers, helps localize where specific task-solving abilities reside in the network.
2. Task dependence of layer importance.
Which layers improve or harm performance when removed is highly task dependent. Sometimes a single layer is critical: for instance, removing layer 25 of Llama-8B on CommonsenseQA causes a 50-point accuracy drop. Removing Llama’s layer 3 improves performance on GSM8K-hard but hurts MATH500; the revers happens when removing layer 11. Removing early layers (1–3) reduces accuracy to near zero on commonsense reasoning tasks (Figure 4), suggesting that certain early layers localize critical task-relevant information.
3. Structured task-specific patterns.
Although pruning is task-specific, related tasks often exhibit similar layer dependencies. Commonsense reasoning tasks (see Figure 4) show importance concentrated in comparable regions of the network. Mathematical reasoning tasks benefit from pruning one to three early layers (e.g., LLaMA layer 3, Mistral layers 6 and 22, Lucie layer 12), but not more (Figures 7, 8, 9). Commonsense and language tasks (ARC, BoolQ, CommonsenseQA, Winogrande, and BIG-Bench) benefit from deleting later layers (Tables 7, 9, 8). This suggests that later layers often play a decoding role for predictions into natural language, which reinforces point 1—pruning them doesn’t harm predictive capability.
We observe stronger pruning gains in reasoning-heavy tasks under zero-shot evaluation. All models showed notable accuracy boosts from one or two layer deletions on mathematical reasoning (e.g., Llama’s large gain on GSM8K-hard, and double-digit gains on MATH500 for Qwen, Mistral, and Lucie). By contrast, knowledge-intensive tasks exhibit more modest improvements (e.g., an 11% gain for Llama on BIG-Bench).
4. Model-specific pruning effects.
Different models display distinct pruning behavior. For example, pruned Lucie achieved a 101% gain on MMLU and double-digit gains on ARC-Challenge, CommonsenseQA, BoolQ and GSM8K-hard. While Qwen-7B, Llama-8B and Mistral share similar architecture and scale, they had modest gains on these data sets. Lucie also benefitted from more substantial pruning than the other models. Interestingly, Lucie was trained on a much smaller dataset (3T tokens vs. 15T for Llama and 13T for Qwen) and also tuned for French conversational proficiency Gouvert et al. (2025), suggesting that some layers for one of our tasks may encode multilingual-specific functionality.
This highlights intriguing interactions between pretraining and pruning efficiency. One hypothesis is that models trained close to their performance ceiling (via large-scale pretraining, instruction tuning or RLHF) yield smaller pruning gains, whereas models trained under limited objectives may benefit more. Notably, even the large corpus trained Qwen-0.5B showed strong pruning efficiency gains.
We experimented with producing pruned models for several tasks, and we saw that a single, general pruning improved speed up without much loss in accuracy across multiple tasks (Appendix E). This shows that judicious pruning can preserve generalization ability across several tasks.
5. Tale is versatile.
Tale can prune: base pre-trained models, instruction-tuned models (as we mainly do here), fine-tuned, and post-trained models with RLHF. Our results show that interleaving pruning and fine-tuning yields compounded benefits that may extend to other training regimes.
8 Conclusions
Tale is a generic toolkit for removing layers not needed for a certain task that yields increased performance and reduced computational costs. We have shown that it can profitably interact with further training or fine tuning, allowing users to further increase their models’ task specific performance. We have also shown that Tale can improve models both at small and larger scales. Tale can benefit high-throughput applications with time constraints–e.g. in multi-agent systems with task-specific agents, real-time recommendation engines, or interactive AI assistants. Tale can also help smaller companies or organizations that face critical trade-offs between model capability and computational efficiency use large language models at scale.
Appendix A Implementation Details
Hardware.
All experiments were conducted on 1 NVIDIA A100 GPU with 80GB memory.
Models.
We applied Tale to five open-weights LLMs of varying scales: Qwen2.5-0.5B-Instruct, Qwen2.5-7B-Instruct, Lucie-7B-Instruct, Mistral-7B-Instruct, and Llama-3.1-8B-Instruct.
Datasets for Tale pruning.
The greedy layer-pruning algorithm was evaluated across nine widely used benchmarks covering reasoning, commonsense, and knowledge-intensive tasks: ARC-Challenge, ARC-Easy, MMLU, Winogrande, GSM8K, MATH500, CommonQA, BIG-Bench, and BoolQ.
Pruning setup.
At each iteration, Tale evaluates all candidate single-layer deletions with respect to validation accuracy. The pruning threshold was defined as the baseline accuracy of the full model, ensuring that pruning never reduces performance relative to the original unpruned model. The iterative procedure terminates once no further layer removals satisfy this criterion.
Fine-tuning setup.
For fine-tuning experiments, we focused on Winogrande and MMLU. We employed LoRA with rank 64, a batch size of 4, and the optimizer paged_adamw_32bit. A cosine learning rate scheduler was used, and models were trained for 10 epochs.
Evaluation.
We measured performance in terms of task accuracy on the test portions of datasets through automatic evaluation, with no human assessment involved, while inference efficiency was reported as relative speedup over the full baseline model. To evaluate MATH500 and GSM8K automatically without any human verification, we force the model to predict the final answer after the reasoning steps in a particular format and then calculate the accuracy. This could lower accuracy from what is expected; but since we used the same evaluation criteria for all the techniques and models and are interested in relative improvements or decreases in performance, the increase/decrease due to our methods from what might have been expected is justified.
Prompting.
For zero-shot and few-shot evaluation, we used task-specific prompts. Below we show the prompt used for datasets, consisting of a system instruction :
ARC-E & ARC-C System Prompt You are a Science expert assistant. Your task is to answer multiple-choice science questions at grade-school level. Each question has four answer choices, labeled A, B, C, and D. For each question: - Carefully read the question and all answer choices. - Select the single best answer from the options (A, B, C, or D). - Respond only with the letter of the correct answer, and nothing else—no explanation or extra words. Be precise and consistent: Only the answer letter.
Bigbench System Prompt ”You are a boolean expression evaluator. You must respond with exactly one word: either ’True’ or ’False’. Do not provide explanations, steps, or any other text. Only respond with ’True’ or ’False’.”
BOOLQ System Prompt ”You are a helpful assistant that answers True/False questions based on given passages. Read the passage carefully and determine if the question can be answered as True or False based on the information in the passage. ”Respond with only ’A’ for True or ’B’ for False.”
CommonQA System Prompt ”You are a helpful assistant that answers multiple-choice questions requiring commonsense knowledge and reasoning. Read each question carefully and select the most logical answer from the given options based on common knowledge and reasoning. Respond with only the letter of your chosen answer (A, B, C, D, or E).”
GSM8K System Prompt ”You are a math problem solver. Solve the given math problem step by step. ” ”Show your complete reasoning and calculations. ” ”At the end, write your final answer after ’####’ like this: #### [your final numerical answer]””
MMLU System Prompt ”You are a helpful assistant that answers multiple-choice questions across various academic subjects including humanities, social sciences, STEM, and professional fields. Read each question carefully and select the best answer from the given options. Respond with only the letter of your chosen answer (A, B, C, or D).”
Winogrande System Prompt You are a careful math problem solver. Show complete step-by-step reasoning and all calculations needed to arrive at the answer. Use clear, numbered or labeled steps so the reasoning is easy to follow. IMPORTANT (formatting): After the full reasoning, write the final answer on a new line by itself in exactly this format: ####
integer
<integer> must be digits only, optionally with a leading “-” for negatives (e.g., -7). Do not add words, punctuation, units, or commentary on the same line as the #### line. The #### line must be the final line of the output (nothing may follow it). Assume all problems expect integer answers; ensure the final line contains a single integer.
Appendix B Number of parameters per layer for each model
Model LLaMA 3.1 8B Qwen 2.5 7B Mistral 7B Lucie 7B Qwen 2.5 0.5B
Parameters 218,112,000 233,057,792 218,112,000 192,946,176 14,912,384
Table 3: Model parameter counts comparison. LLaMA 3.1 8B, Mistral 7B and Lucie 7B has 32 layers, Qwen 2.5 7B has 28 layers and Qwen 2.5 0.5B has 24 layers.
Appendix C Intution behind Tale
(a) ARC-Challenge
(b) MMLU
Figure 5: Layer-wise output performance for LLaMA models: results when generating predictions from intermediate layers 1 through 32 on three different datasets.
Appendix D Results on few-shot learning on larger models
Dataset Lucie 7B few-shots
Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 69.2 72.36 9 1.41 71.27 12 1.68
ARC-Challenge 49.31 55,17 9 1.39 51.72 13 1.67
BoolQ 77.6 79.10 6 1.22 78.5 10 1.27
MMLU 41.02 43.44 7 1.26 41.48 11 1.55
COMMONQA 55.4 69.7 3 1.22 57.10 17 2.02
WINOGRANDE 52.8 56.90 12 1.58 53.30 17 1.74
BIG-Bench 68.8 77.20 9 1.61 72 15 2.23
GSM8K-HARD 26.97 29.21 1 1.03 26.97 2 1.1
Table 4: Results of Lucie 7B across nine benchmarks. All tested on 5-shots, except gms8k on 8-shots Performance (%) cells are color-coded: green = gain, red = decline, and gray = near-neutral change compared to baseline.
Dataset LLaMA 3.1 8B few-shots
Baseline Best Model BSBA
Perf. Perf. #D Sp. Perf. #D Sp.
ARC-Easy 90.36 92.18 2.01% $\uparrow$ 4 1.14 90.91 8 1.37
ARC-Challenge 78.2 83.10 6.27% $\uparrow$ 3 1.17 78.62 9 1.42
BoolQ 82.7 85.3 3.1% $\uparrow$ 4 1.11 83.0 6 1.22
MMLU 59.2 62.38 5.37% $\uparrow$ 4 1.14 59.57 7 1.26
COMMONQA 73.30 75.30 2.72% $\uparrow$ 6 1.22 73.80 7 1.32
WINOGRANDE 57.01 60.1 5,26% $\uparrow$ 3 1.1 57.02 8 1.3
BIG-Bench 70.0 83.60 19,43% $\uparrow$ 5 1.2 81.20 15 1.83
GSM8K-HARD 60.67 60.67 0 1 60.67 0 1
MATH500 44.00 49.00 11.36% $\uparrow$ 1 1.02 45.00 2 1.03
Table 5: Results of LLaMA 3.1 8B across nine benchmarks. All tested on 5-shots, except gms8k and MATH500 on 8-shots
Appendix E A tunable metric for finding accuracy vs. speed up optimization
To systematically select among these candidates according to user priorities, we propose the Accuracy–Efficiency Harmonic Mean (AE-HM):
$r_{A}=\frac{\text{Acc}(\text{Model})}{\text{Acc}(\text{Baseline})},\qquad\text{AE-HM}(\text{Model})=\frac{(1+\lambda^{2})r_{A}\,S}{\lambda^{2}S+r_{A}}=\frac{1+\lambda^{2}}{\frac{\lambda^{2}}{r_{A}}+\frac{1}{S}}$ (2)
where $S$ denotes the relative inference speedup and $\lambda$ controls the relative importance of accuracy versus efficiency. The user can set AE-HE’s parameter $\lambda$ to desired specifications: if $\lambda>1$, we prioritize $r_{A}$; if $\lambda<1$ we prioritize Speedup.
By computing AE-HM for candidate models, we can automatically identify the model with the highest score for a given task or a set of tasks given a particular AE-HM parameter setting:
$M_{\text{best-compromise}}=\arg\max_{i}\text{AE-HM}(M_{i})$ (3)
Figure 6: Relative Gain comparison across datasets. LLaMA $\beta=3$
Appendix F Deleted Layers in each Model and Benchmark
Dataset Best Model BSBA
ARC-Easy 19 25 27 28 19 20 21 24 25 26 27 28
ARC-Challenge 19 22 27 19 20 21 22 23 24 26 27 28
BoolQ 19 25 26 32 15 19 21 22 25 26 30 32
MMLU 20 21 27 28 20 21 22 24 27 28 32
CommonQA 21 22 27 28 31 32 21 22 23 27 28 31 32
Winogrande 20 22 24 17 19 20 22 24 26 29 32
BIG-Bench 11 16 20 21 26 10 11 16 20 21 22 23 24 26 27 28 29 30 31 32
MATH500 28 24 28
Table 6: Deleted layers represented as color-coded inline numbers. Blue = Best Model, Orange = BSBA for LlaMA 3.1 8B with few-shots.
Table 7 shows how using AE-HM allows us to bring model size down effectively on our BSBA Llama model with 0 shot performance on our nine data sets. The BSBA LLama model had speed up gains between 27 and 46% on our various benchmarks and maintained performance at or above original model levels (See Table 7).
Dataset Best Model BSBA
ARC-Easy 19 20 21 29 32 19 20 21 22 25 27 29 32
ARC-Challenge 19 20 23 27 19 20 21 23 25 27 28
BoolQ 21 23 28 18 21 22 27 28 32
MMLU 21 19 21 22 24 25 26 27 28 31
CommonQA 19 23 28 19 22 23 26 27 28
Winogrande 23 24 26 32 20 21 22 23 24 25 26 27 29 31 32
BIG-Bench 14 20 22 28 29 14 18 20 21 22 23 24 28 29 31 32
GSM8K-Hard 3 3 21 22 25 26 27 29
Table 7: Deleted layers represented as color-coded inline numbers. Blue = Best Model, Orange = BSBA for LlaMA 3.1 8B 0 shot.
Dataset Best Model BSBA
ARC-Easy 19 22 28 6 19 22 24 26 27 28
ARC-Challenge 27 28 7 22 23 26 27 28
BoolQ 18 21 27 28 12 19 21 22 26 27 28
MMLU 22 23 26 27 28 18 22 23 26 27 28
CommonQA 22 28 6 21 22 23 27 28
Winogrande 22 26 27 6 20 22 25 26 27
BIG-Bench 10 19 23 25 26 27 10 19 23 25 26 27
Table 8: Deleted layers represented as color-coded inline numbers. Blue = Best Model, Orange = BSBA for Qwen 2.5 7B zero-shot.
Dataset Best Model BSBA
ARC-Easy 15 16 23 24 27 28 13 15 16 18 19 20 21 22 23 24 25 27 28
ARC-Challenge 16 18 20 21 23 25 26 15 16 18 19 20 21 22 23 25 26 28
BoolQ 8 17 25 28 29 5 8 11 12 13 14 15 16 17 19 20 23 25 26 27 28 29 31
MMLU 11 12 15 16 20 21 22 28 5 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 28 30 31
CommonQA 11 12 27 11 12 13 15 16 17 18 19 20 21 22 23 24 25 27 28
BIG-Bench 6 7 15 17 20 21 25 26 27 6 7 13 15 17 19 20 21 22 24 25 26 27 28 29
GSM8K-Hard 12 12 21 23
Table 9: Deleted layers represented as color-coded inline numbers. Blue = Best Model, Orange = BSBA for Lucie 7B 0 shots.
Dataset Best Model BSBA
ARC-Easy 21 22 24 26 29 21 22 23 24 25 26 29 30 32
ARC-Challenge 22 24 25 27 28 30 21 22 24 25 26 27 28 30
BoolQ 17 22 23 24 27 32 12 17 21 23 24 25 27 28 32
MMLU 24 30 22 23 24 25 26 27 30 32
CommonQA 19 22 25 28 19 21 22 24 25 28 32
Winogrande 18 19 20 22 23 24 26 27 31 32 4 13 18 19 20 22 23 24 26 27 29 31 32
BIG-Bench 3 5 15 22 23 24 26 27 28 3 5 14 15 18 22 23 24 26 27 28
GSM8K-Hard 6 22 6 11 22 28
Table 10: Deleted layers represented as color-ccdinline numbers. Blue = Best Model, Orange = BSBA for Mistral zero-shot.
Appendix G Common pruned layers model
Group Dataset Baseline Pruned Model speedup
Common-sense ARC-Easy 87.0 87.82 1.2
ARC-Challenge 75.86 75.00 1.21
CommonQA 72.20 64.70 1.1
Winogrande 54.20 50.57 1.13
Reading BoolQ 85.0 85.5 1.17
BIG-Bench 75.2 67.2 1.1
Table 11: Accuracy of LLaMA-3.1-8B (baseline) versus a pruned variant obtained by dropping layers selected through BSBA. For each task, BSBA identified removable layers, and we retained the intersection of layers that appeared in at least 75% of tasks within the Common-sense group (layers 19, 22, 23, 27) and (layers 18, 21, 22, 28, 32) for Reading Comprehension tasks. These layers were then pruned globally from the model, and performance was re-evaluated across tasks. Speedup is reported relative to the baseline.

### v2

@@
 Abstract
-In this paper we introduce Tale, Task-Aware Layer Elimination, an inference-time algorithm that prunes entire transformer layers in an LLM by directly optimizing task-specific validation performance.
-We evaluate Tale on 9 tasks and 5 models, including LLaMA 3.1 8B, Qwen 2.5 7B, Qwen 2.5 0.5B, Mistral 7B, and Lucie 7B, under both zero-shot and few-shot settings.
-Unlike prior approaches, Tale requires no retraining and consistently improves accuracy while reducing computational cost across all benchmarks.
-Furthermore, applying Tale during finetuning leads to additional performance gains.
-Finally, Tale provides flexible user control over trade-offs between accuracy and efficiency.
-Mutual information analysis shows that certain layers act as bottlenecks, degrading task-relevant representations.
-Tale’s selective layer removal remedies this problem, producing smaller, faster, and more accurate models that are also faster to fine-tune while offering new insights into transformer interpretability.
+Large Language Models (LLMs) are typically deployed using a fixed architecture, despite growing evidence that not all layers contribute equally to every downstream task.
+In this work, we introduce TALE (Task-Aware Layer Elimination), an inference-time method that improves task performance by selectively removing layers that are irrelevant or detrimental for a given task.
+TALE optimizes task-specific validation performance, yielding a task-adapted architecture without retraining or modifying model weights.
+Across 9 tasks and 5 model families, under both zero-shot and few-shot settings, we show that TALE consistently matches or surpasses baseline performance while simultaneously reducing computational cost, outperforming general and layer-wise pruning approaches such as SLEB.
+Beyond inference-time gains, TALE synergizes with fine-tuning and few-shot learning, where task-adapted architectures lead to additional performance improvements.
+Computing TALE for a new task requires modest resources (1–2 GPU hours on an A100), making it a practical and deployable solution for task-specialized LLM inference.
+TELL-TALE: Task Efficient LLMs with Task Aware Layer Elimination
+Omar Naim1thanks: Equal contribution.   Krish Sharma [^1] footnotemark: 1   Niyar R Barman 1   Nicholas Asher  [^1] footnotemark: 1 1 IRIT, Toulouse , France
 1 Introduction
-While Large Language Models (LLMs) have achieved remarkable success, they come with large computational overheads.
-This may prevent users in resource limited organizations or with high-throughput applications from using more capable models.
-This limitation has spawned an important research area on model compression.
-We focus here on the branch known as model pruning.
-Although it often reduces computational overhead, pruning typically requires complex implementation procedures, extensive retraining or fine-tuning, and may result in unpredictable performance degradation (Section 2 details related work).
-In contrast, having observed that not all layers contribute to a particular task, we propose a lightweight, greedy, iterative layer pruning algorithm, Task Aware Layer Elimination (Tale).
-Tale operates at inference time, is hardware agnostic, directly optimizes for task-specific accuracy at each pruning step and consistently offers improved results over the original model.
-This improvement persists in interactions with fine tuning on our tasks.
-As illustrated in Figure 1 and detailed in Section 3, Tale systematically evaluates all possible single-layer removals at each iteration, selecting the layer whose elimination results in the highest validation accuracy.
-This process continues iteratively until performance improvements fall below a predefined threshold, ensuring that only layers with minimal or negative impact on task performance are removed.
-Tale improves task specific accuracy and provides moderate computational reductions with minimal implementation complexity.
-This can be extremely useful for agentic systems with LLMs for dedicated tasks; TALE will make those LLMs perform better and more quickly.
-Our layer-wise pruning leverages the modular nature of transformer architectures, where each layer performs a complete transformation of the input representation through attention and feedforward mechanisms.
-This architectural property enables the removal of entire layers without requiring modifications to the remaining network structure.
-We evaluate Tale with five LLMs, LLaMA 3.1 8B, Qwen 2.5 7B, Qwen 2.5 0.5B, Mistral 7B and Lucie 7B, on 9 diverse benchmark datasets (Sections 4 and 5) both in zero-shot and few-shot settings.
-We also investigated interactions between pruning with Tale finetuning.
-We find consistent improvements in both accuracy and computational efficiency in all cases.
-This suggests pre-trained LLMs contain redundant or even detrimental layers for a given downstream task.
-We analyze layer flow using the notion of mutual information (MI) to support this hypothesis (Section 6).
-Additionally, our experiments show Tale’s potential as a tool for understanding layer function in and across models (Section 7), thus aiding model interpretability.
+The substantial computational costs of Large Language Models (LLMs) can prevent resource-constrained organizations and those with high-throughout applications from leveraging more capable models.
+This has led to a search for methods that boost task-specific performance and reduce computation costs.
+The use of multi-agent systems, with LLMs specialized for a particular role, has intensified the need for such methods.
+Yet finding such methods has proved elusive.
+Fine-tuning can increase task performance but does not reduce inference costs and requires significant training overhead and data.
+General pruning reduces computation costs but typically demands significant retraining and often results in substantial performance degradation on downstream tasks.
+Tale, Task Aware Layer Elimination addresses this need.
+Tale increases task performance while modestly reducing computational overhead.
+Tale evaluates all possible single-layer removals at each iteration, selecting the layer whose elimination results in the highest validation accuracy.
+This process repeats until performance falls below a user predefined threshold.
+Tale is a simple, hardware agnostic, greedy algorithm requiring no retraining that runs at inference time.
+Peer et al.
+(2022)’s similar greedy approach, on the other hand, requires retraining.
+Unlike all pruning methods of which we are aware, Tale optimizes for task-specific accuracy at each pruning step.
+We believe this is why Tale significantly improves results over the original model and interacts synergistically with fine-tuning.
+Layer analysis on earlier models observed that not all layers contribute equally to a given taskDe Vries et al.
+(2020); Dalvi et al.
+(2020); Sajjad et al.
+(2023).
+Our study of state of the art, middle-sized transformers sharpens these observations: the layers most or least useful for a given LLM are highly dependent on the target task.
+These findings validate the attractiveness of Tale’s simple, task-specific, accuracy-driven design.
+We showcase Tale ’s gains on five modern LLMs, LLaMA 3.1 8B, Qwen 2.5 7B, Qwen 2.5 0.5B, Mistral 7B and Lucie 7B, with 9 diverse benchmark datasets (Sections 4 and 5).
+We also compare Tale to previous training-free pruning methods on two older models (Llama 2 7B and 13B) and show that Tale achieves substantially higher accuracy with efficiency gains.
+With these benchmarks and model families, we show that removing task-misaligned layers can reduce inference cost and, in many cases, improve accuracy.
+Importantly, these improvements arise without retraining, architectural modification, or changes to the underlying model weights, and persist under alternative data splits, random seeds, and evaluation protocols.
+TALE complements fine-tuning and applies post hoc to existing checkpoints, making it a practical tool for deploying task-specialized LLMs.
+Figure 1: How TALE works to produce BEST and Best Speedup above Baseline (BSBA) models.
 2 Related work
-Zhu et al.
-(2024) distinguishes four primary approaches to model compression: model pruning, quantization, low-rank approximation, and knowledge distillation.
-Our work focuses on pruning, which can be categorized into unstructured, structured, and semi-structured methods.
-Unstructured pruning removes individual parameters, resulting in irregular, sparse structures Han et al.
-(2015b); Chen et al.
-(2015); Srinivas & Babu (2015); structured pruning eliminates entire components such as neurons, attention heads, or layers while maintaining the overall network structure He et al.
+While there is considerable work on model speedup through various pruning methods (pruning sparsity, acceleration, model compression), very few methods concentrate on boosting accuracy and almost no work has used compression methods to improve performance, which is what Tale does.
+Structured pruning, which eliminates entire components such as neurons, attention heads, or layers He et al.
 (2017); Voita et al.
 (2019); Lagunas et al.
-(2021); Men et al.
-(2024).
-Semi-structured pruning combines fine-grained control with structural regularity, and has been explored in recent work Li et al.
-(2023); Frantar & Alistarh (2023b); Sun et al.
-(2024).
-Early pruning methods leveraged second-order information for structured pruning LeCun et al.
-(1989); Hassibi et al.
-(1993), but the field has since shifted toward computationally simpler, magnitude-based approaches that prune parameters by importance scores Han et al.
-(2015a); See et al.
-(2016); Narang et al.
-(2017)
-Model pruning has benefited from information-theory.
-Building on the information bottleneck view of deep networks, Tishby et al.
-(2000); Tishby & Zaslavsky (2015) conceptualize layers as filters that preserve task-relevant information while discarding irrelevant input.
-This has led to model pruning proposals based on mutual information (MI).
-Fan et al.
-(2021) propose a layer-wise strategy that leverages MI estimates to reduce hidden dimensionality in a top-down manner.
-MI-guided approaches demonstrate that information-theoretic criteria can balance representation quality with model compression (Ganesh et al., 2020; Westphal et al., 2024).
-A central challenge, however, is the difficulty of estimating MI.
-Neural estimators such as MINE Ishmael Belghazi et al.
-(2018) provide principled solutions for high-dimensional settings, but in practice, probing classifiers Belinkov (2022) remain the dominant tool due to their efficiency and interpretability.
-With regards to large transformers, Zhang & Papyan (2025) proposes a pruning strategy based on matrix approximations.
-Xia et al.
-(2023) demonstrates that pruning 7B models into 1.3B and 2.7B variants via structured layer and hidden-dimension pruning, combined with dynamic batch loading for fine-tuning, can yield submodels outperforming comparably sized models trained from scratch, though not matching the original model.
-Complementary work Frantar & Alistarh (2023a); Zhang et al.
-shows that contiguous blocks, especially attention layers, can often be pruned with limited performance loss, although critical “cornerstone” layers Zhang et al.
-remain essential for reasoning.
-Kim et al.
-(2024) investigates lock-level pruning based on weight importance, but these methods typically require fine-tuning techniques to recover baseline accuracy.
-Most pruning methods compress models while attempting to match the original performance but often suffer accuracy degradation, which requires retraining or fine-tuning Xia et al.
-(2024).
-Improvements are generally measured relative to small models rather than original unpruned baselines.
-To the best of our knowledge, Tale is the only pruning method reported to improve over the original model’s performance without requiring any additional training or fine-tuning.
-Figure 1: Illustration of Tale layer elimination.
-Candidate layers (yellow) are tested for removal, and the best-performing ones above the threshold are permanently dropped (red) until no further improvement is possible.
-3 Tale , our Greedy-Selection Algorithm
-3.1 Basics and Intuitions
+(2021) is relevant to Tale .
+Frantar and Alistarh (2023a); Zhang et al.
+(2024) prune contiguous blocks with minimal performance loss.
+Similar to earlier methods that use representation similarity Dalvi et al.
+(2020), SLEB Song et al.
+(2024) removes entire layers based on the cosine similarity of their representations, which is the closest layer-wise baseline to Tale .
+SLEB’s layer selection is validated using a general metric (perplexity) on a linguistic dataset.
+SliceGPT Ashkboos et al.
+(2024) prunes layer dimensions via Principal Component Analysis.
+However, representational similarity, whether measured by cosine similarity, Center Kernel Alignment (Dalvi et al., 2020), or variance under PCA (Sajjad et al., 2023) is not a good measure for assessing a layer’s importance for a task.
+In Section 5.1 we show that representational similarity and validation with perplexity are sub-optimal for maximizing performance on specific downstream tasks.
+SparseGPT Frantar and Alistarh (2023b) and Wanda Sun et al.
+(2023) perform weight-level unstructured pruning based on local reconstruction criteria or magnitude-activation products, achieving high sparsity but often degrading task-specific reasoning and linguistic abilities.
+Methods like OWL(Yin et al., 2024) and FLAP (An et al., 2024) use non uniform sparsity pruning to achieve important efficency gains, while Tang et al.
+(2025) uses an evolutionary algorithm and a fitness criterion on a selection of data to find models with a given sparsity.
+None of these approaches provide real improvements on accuracy for standard benchmarks, and for the most part they show significantly decreased performance from the baseline.
+Peer et al.
+(2022) use a preset accuracy goal and a greedy algorithm like ours and show that their greedy-layer pruning (GLP) outperforms Top-layer pruning Dalvi et al.
+(2020) in almost all cases.
+Although superficially similar, GLP is very different from Tale.
+First, GLP retrains the pruned model on a downstream task T and compare the retrained model’s results on T to the those of the baseline model.
+Tale, on the other hand, involves no retraining.
+Even with retraining GLP’s reported accuracies for GLP models are far below our best models.
+Second, GLP restricts its algorithm to remove an a priori number of layers irrespective of optimization.
+Tale selects the number of layers that optimizes performance without setting a number a priori of layers to remove.
+Tale differs from other training free methods in granularity and optimization objective: Tale is a structured, layer-level method that is explicitly optimized for task-specific accuracy, whereas the aforementioned training-free baselines are designed for compression using predefined limits and linguistic or reconstruction criteria.
+Section 5 compares various approaches with Tale , showing that Tale consistently and substantially improves accuracy over the unpruned baseline, a benefit not reliably achieved by other methods.
+3 Basics and Intuitions
 A transformer maps a sequence of input vectors $(x_{1},\cdots,x_{n})$ to a corresponding sequence of output vectors through a stack of L layers.
@@
 Removing layer $\ell$ from this pipeline simply redirects the flow such that $X^{(\ell-1)}\rightarrow X^{(\ell+1)}$, a property that makes the architecture naturally amenable to layer-wise pruning.
-Our initial intuition for Tale came from examining the behavior of partial forward passes.
+Our intuition for Tale came from examining the behavior of partial forward passes.
 Let $h^{(k)}$ denote the hidden representation after $k$ layers.
@@
 $\hat{y}^{(k)}=\mathrm{softmax}(W_{\text{out}}h^{(k)}).$
-We then compared the performance of $\hat{y}^{(k)}$ across different values of $k$.
-Surprisingly, we observed that for many tasks, intermediate layers ($k<\text{L}$) achieved higher accuracy than the final layer L (Figure 5).
-This indicated that additional depth does not always translate into better task-specific performance: some layers contribute marginally, while others introduce representational noise.
-This experiment motivated our central hypothesis: not all layers in an LLM are equally useful, and selectively removing redundant layers can preserve—or even improve—downstream accuracy.
-Tale (Task-Aware Layer Elimination) formalizes this intuition into a principled, iterative pruning strategy.
-3.2 Tale
-Tale is a greedy iterative layer pruning algorithm for pre-trained open-weights LLM compression that systematically removes layers while preserving or even improving model performance (Algorithm 6).
-Starting with a full pre-trained model, Tale evaluates all possible single-layer removals at each iteration, computing the validation accuracy for each candidate pruned architecture.
-The layer whose removal results in the highest accuracy is permanently eliminated from the model, and this compressed architecture becomes the baseline for the next iteration.
-This process continues iteratively until the performance improvement falls below a predefined threshold, at which point the algorithm terminates and returns the most compressed model that maintains performance above the specified threshold.
-Our approach directly optimizes for task-specific accuracy at each pruning step, ensuring that only layers with minimal impact on the target objective are removed.
-This exhaustive evaluation strategy, while computationally intensive during the pruning phase, provides strong empirical guarantees about the optimality of each pruning decision within the greedy framework.
-Algorithm 1 Tale : Greedy Iterative Layer Pruning
+When comparing the performance of $\hat{y}^{(k)}$ across different values of $k$, we surprisingly observed that for many tasks, intermediate layers ($k<\text{L}$) achieved higher accuracy than the final layer L (Figure 7).
+Thus, additional layers did not always improve task-specific performance: some layers contribute marginally, while others introduce representational noise.
+Not all layers in an LLM are equally useful, and selectively removing redundant layers can preserve—or even improve—downstream accuracy.
+Tale (Task-Aware Layer Elimination) formalizes this intuition into a principled, iterative optimization strategy.
+Figure 2: Accuracy progression of Tale across 3 benchmark datasets for LLaMA 3.1 8B.
+Each curve represents the accuracy at successive iterations.
+The $\star$ denotes the best-performing layer drop configuration, while the $\bullet$ highlights the Best Speed up with at least Baseline Accuracy (BSBA) configuration.
+Plots for all tasks are in Appendix E.
+Algorithm 1 Tale : Iterative Layer Pruning
 1:Pre-trained model $\mathcal{M}$ with $L$ layers; validation set $\mathcal{D}_{val}$; performance threshold $\epsilon$
@@
 16:return $\mathcal{M}^{*}$
+3.1 Tale
+Tale is a greedy, iterative layer pruning algorithm for pre-trained open-weights LLM compression that systematically removes layers while preserving or even improving model performance (Algorithm 1).
+Starting with a full pre-trained model, Tale evaluates all possible single-layer removals at each iteration, computing the validation accuracy for each candidate pruned architecture.
+The layer whose removal results in the highest accuracy is permanently eliminated from the model, and this compressed architecture becomes the baseline for the next iteration.
+This process continues iteratively until the performance improvement falls below a predefined threshold.
+For this paper we take baseline task accuracy - 8% to be the threshold, which allows the accuracy during our search to go slightly below baseline to capture eventual, substantial increases.
+We have found no cases where the trajectory later goes above the baseline.
+Once the threshold is reached, the algorithm terminates and returns the most compressed model with performance above the threshold.
+Task performance optimization trajectories Figure 2 visualizes three iterative model optimization processes for LLaMA 3.1 8B with Tale .
+Each curve tracks accuracy as layers are progressively removed.
+The graphs reveal a general pattern across almost all our tasks: the first iteration of TALE typically provides a large boost in accuracy, followed by slight increases or decreases; they then follow a monotonic decreasing path to accuracies below the baseline and eventually to 0.
+A full set of curves is in Figure 11.
+The curves in themselves are worthy of future study.
+Three consistent patterns emerge from these trajectories: (i) Tale identifies compressed BEST models that outperform the original across diverse tasks, with $\star$ markers lying strictly above baseline.
+Best Speed up with at least Baseline Accuracy (BSBA) are the models with best speed up above Tale’s stopping point (with $\bullet$).
+(ii) Accuracy improvements persist across multiple pruning steps before diminishing returns, showing that substantial redundancy exists even in carefully tuned pretrained models.
+(iii) Pruning dynamics are task-specific: datasets such as ARC-Easy and MMLU tolerate deeper pruning while continuing to improve, whereas reasoning-heavy tasks like GSM8K-Hard converge earlier, reflecting heterogeneous layer importance across domains.
 4 Benchmarks and Datasets
 We evaluate Tale across a diverse suite of nine benchmarks spanning reasoning, language understanding, and commonsense knowledge.
-For mathematical reasoning, we include GSM8K-Hard, a curated subset of GSM8K Cobbe et al.
+For mathematical reasoning, we use GSM8K-Hard, a curated subset of GSM8K Cobbe et al.
 (2021) with more than five premises per question to increase difficulty, and MATH500 Hendrycks et al.
-(2021b), a benchmark for symbolic and arithmetic reasoning (for evaluation details see Appendix A).
-For language understanding, we consider MMLU Hendrycks et al.
+(2021b).
+(for evaluation details see Appendix A).
+For language understanding commonsense reasoning and multi-task generalization, we use MMLU Hendrycks et al.
 (2021a) and BoolQ Clark et al.
-(2019), while Winogrande Sakaguchi et al.
+(2019); Winogrande Sakaguchi et al.
 (2021), CommonsenseQA Talmor et al.
 (2019), andBIG-Bench Srivastava et al.
-(2023) capture commonsense and multi-task generalization.
-Finally, we include both ARC-Easy and ARC-Challenge Clark et al.
-(2018), which evaluate scientific and factual reasoning at varying difficulty levels.
-Together, these nine datasets cover a broad spectrum of downstream challenges and allow us to assess both the generality and task-specific benefits of our pruning strategy.
+(2023), ARC-Easy and ARC-Challenge Clark et al.
+(2018).
+In our study Tale requires only modestly-sized sets for task-specific optimization, ranging from 500 to 1500 examples.
+As seen in Table 3 (Appendix C), once the validation set size exceeds 500 examples, the set of layers dropped stabilizes across all tasks.
+For each benchmark, we partition the available data into three disjoint subsets: (i) a training pool, (ii) a held-out optimization split drawn from the training pool, and (iii) a final evaluation split (the official test set when available, or a held-out validation set otherwise).
+TALE uses only the optimization split to select which layers to remove for a given task.
+The evaluation split is never used during layer selection.
+All results reported throughout the paper correspond to performance on this unseen evaluation split.
+When official test sets are unavailable, we follow prior work and report results on held-out validation splits that are disjoint from the data used for layer optimization.
+To see results with two different training subsets see Tables 2 and 17.
 5 Results
-We evaluate Tale across five medium-scale models (LLaMA 3.1 8B, Mistral 7B, Lucie 7B, Qwen 2.5 7B) and one smaller model (Qwen 2.5 0.5B), spanning nine benchmarks that cover commonsense reasoning, reading comprehension, and mathematical problem solving.
-All experiments are conducted in the zero-shot setting unless otherwise noted.
-[^1] Code available at https://anonymous.4open.science/r/tale/ Table 3 summarizes model configurations.
+We emphasize that all TALE pruning decisions are made using held-out optimization splits drawn from the training data only.
+Test (or final validation) sets are never used for layer selection and are accessed exactly once for evaluation.
+This protocol ensures that reported gains reflect genuine generalization rather than selection bias.
+5.1 Comparisons
+We first compare Tale ’s best models to state of the art alternatives on two older models. We then analyze our method’s robustness on five more modern models.
+Most of the state of the art methods for training-free pruning used Llama2-7b and Llama2-13b with LM Eval evaluation (accuracy).
+Figure 3 shows that state of the art methods never surpassed the baseline accuracies on the tasks we measured.
+On the other hand Tale produced best models that consistently surpassed baseline accuracies on the task.
+Figure 3: Performance comparison of pruning methods on LLaMA-2 models.
+Evaluation of seven pruning approaches using LM-Eval accuracy across four zero-shot benchmarks for LLaMA-2-7B (left) and LLaMA-2-13B (right).
+Methods are shown in a blue gradient from light (Baseline) to dark (Tale).
+Tale achieves the highest accuracy across all tasks while maintaining equivalent sparsity to SLEB, which performed second best to Tale.
+Method outperformed all structured pruning methods (SpareGPT, Wanda, SliceGPT) on both model sizes, demonstrating effective layer dropping without accuracy loss.
+Prior work on structured layer pruning, most notably SLEB, has already evaluated naive depth-based baselines such as removing the top-$k$ or bottom-$k$ transformer layers and demonstrated that these strategies are consistently inferior to task- or redundancy-aware pruning.
+Since our focus is on task-aware layer selection rather than re-criticizing uniform depth trimming, we do not repeat these experiments here but compare Tale directly against SLEB and other training-free pruning methods that represent the strongest existing baselines.
+Why were methods geared towards tasks so inferior to Tale?
+Tale is focused on optimizing accuracy on generated output or with LM Eval.
+Other methods used representational similarity to find redundant layers.
+So we used a representational similarity technique (cosine similarity or cossim) to guide Tale.
+Tale registered drops similar to those seen with SLEB and other similarity driven optimization methods.
+On ARC-Easy, for instance, cossim led TALE to drop 2 layers, dropping task accuracy from Llama’s baseline of 79.5 to a pruned model accuracy of 58.5 with a time speed up of 1.32.
+Table 16 in the Appendix provides a direct comparison with SLEB, a similarity based method.
+We conclude that similarity based optimization for task performance is not competitive with Tale .
+As SLEB and other methods use perplexity as a check on pruning, we also used Tale to optimize models for perplexity.
+Tale provided minimal increases in perplexity with speed up (see Table 15.
+More importantly, optimizing for perplexity did not translate into better performance on downstream tasks, contra Song et al.
+(2024).
+5.2 Robustness
+We next evaluated Tale on our benchmarks in 0-shot settings across four medium-scale models (LLaMA 3.1 8B, Mistral 7B, Lucie 7B, Qwen 2.5 7B) and one smaller model (Qwen 2.5 0.5B).
+We wanted to test Tale with respect to improvements in accuracy on actual generated output.
+All base models could obey this prompt.
+Figure 4 compares accuracy of four mid-sized baseline models against pruned counterparts selected using validation splits drawn from the training data; all reported results are evaluated on a disjoint test (or held-out validation) set.
+Across all benchmarks and base models, Best models yield consistent accuracy gains though percentage gains vary across models and tasks.
+[^1] Code available at https://anonymous.4open.science/r/tale/ On Arc Challenge, Llama Best has the lowest increase in accuracy (though still significant) at 1.6%, while Qwen 7b gains 6.3%.
+Reasoning tasks (Math500, GSM8K) benefited most from Tale with gains ranging from 23% to 51% across all models.
+Similar results on the larger models using LM-Eval are in the appendix.
+The fact that we see the same behavior under two different evaluations is evidence that Tale is capturing a real phenomenon, not an artifact of evaluation.
+Figure 5 shows that shifts due to different seeds were minimal.
+Table 2 in the appendix shows numerical values for the averaged results of Tale with 5 different seeds along with variance on both Llama, Qwen, Lucie and Mistral best models.
+Given the low variance, we take this to show that Tale is robust.
+Figure 4: Zero-shot performance comparison across four language models with layer dropping.
+Light blue bars show baseline performance, while dark blue bars represent best model performance after strategic layer dropping.
+Circled numbers indicate the number of layers dropped (#D) for each task.
+Results demonstrate that dropping 1–10 layers can improve or maintain performance across most benchmarks, with notable gains on mathematical reasoning tasks (e.g., LLaMA 3.1 8B: 39.0% → 59.0% on GSM8K-HARD with only 1 layer dropped).
+This suggests significant redundancy in deeper language model architectures.
+Figure 5: Model Stability Across Datasets.
+Heatmap visualizing the standard deviation ($\sigma$) of the final accuracy scores for the Best Model variant of four different Large Language Models (LLMs) across nine datasets.
+The results are aggregated from five independent runs (seeds).
+Lighter green indicates higher stability (lower $\sigma$); darker green, lower stability (higher $\sigma$).
+LLaMA 3.1 8B shows the overall lowest variance.
+5.3 Computational Cost and Amortized Efficiency
+Tale requires a modest, one-time computation to identify the optimal layer-set for a given task, which is then amortized over the model’s entire inference lifetime.
+For an $L$-layer model and a validation set of size $V$, the time of pruning process is proportional to $O(I\cdot L\cdot V\cdot T_{\text{layer}})$, where $I$ is the number of pruning iterations (details in Appendix D).
+For LLaMA 3.1 8B across our benchmarks ($L=32$, $V\approx 500-1500$), the pruning required approximately 1 to 2 GPU-hours on a single A100.
+Takeaways.
+Tale consistently uncovers high accuracy and high accuracy/high efficiency models.
+By balancing task fidelity with computational savings, it enables both accuracy-focused and efficiency-focused deployment.
+[^2] Tale comes with a tunable selection metric for choosing best candidate trade-offs in Appendix J.
+Even strong, larger models like Qwen 7B see significant improvements, but so do small models (Qwen 0.5B).
+The stability of the scores under different seeds and different training splits (see Appendix Table 2) underscores the robustness of Tale and the value of its task directed optimization, rather than one-size-fits-all pruning heuristics for model optimization.
+5.4 Interactions between Tale and few shot learning and Fine-tuning
+Few-shot setting.
+As few-shot prompting improves baselines on many tasks,333in particular reasoning tasks like GSM8K and Math500 we tested on Lucie and LLaMA models whether Tale could synergize with few shot prompting to bring higher gains.
+(Appendix Tables 67).
+Tale-pruned variants still achieve higher accuracy in nearly all settings.
+This shows that Tale-induced improvements are largely complementary to gains from in-context learning.
+Fine tuning We also investigated whether Tale’s removing- layers before or after fine-tuning affects model performance.
+Removing layers reduces representational capacity even for Qlora fine-tuing, and so might limit downstream fine-tuning performance compared to baseline instruct-tuned models.
+Our experiments, however, show the opposite: Tale not only preserves fine-tuning results but can even improve accuracy and efficiency.
+We explored four settings: (i) fine-tuning the base model (FT), (ii) applying Tale after fine-tuning (FT $\rightarrow$ Tale ), (iii) pruning first and then fine-tuning (Tale $\rightarrow$ FT), and (iv) pruning first, then fine-tuning, and finally pruning again (Tale $\rightarrow$ FT $\rightarrow$ Tale ).
+Across various benchmarks, we consistently observed mostly moderate and sometimes significant gains after iterating pruning and fine-tuning, especially on Winogrande and GSM8K (Table 1).
+This suggests that pruning can act as a regularizer, simplifying the optimization landscape by removing redundant layers.
+Tale also reduced computation costs for fine-tuning.
+For example, pruning LLaMA-3.1 8B before fine-tuning reduced fine-tuning time by 2–2.5 GPU hours on an A100 (an 18.5% reduction) while simultaneously improving Winogrande performance by $2.4\%$.
+Iteratively applying pruning and fine-tuning allowed us to prune up to 8 layers achieving still higher accuracy (87.37%) than the full fine-tuned model (85.00%).
+Similarly, pruning the fully fine-tuned model yielded a 7-layer reduction while maintaining strong accuracy (86.66%).
+Model Dataset Baseline Tale FT Only Tale $\rightarrow$ FT FT $\rightarrow$ Tale (Tale $\rightarrow$ FT) $\rightarrow$ Tale
+Perf. #D Perf. #D Perf. #D Perf. #D Perf. #D Perf. #D
+Llama 3.1 8B Winogrande 53.83 0 56.67 4 85.00 0 87.06 4 86.74 7 87.37 8
+MMLU 54.87 0 59.90 1 63.62 0 63.49 1 64.21 2 64.01 2
+CommonQA 72.20 0 75.30 3 81.88 0 81.80 3 83.40 3 82.90 6
+GSM8K 15.07 0 37.08 3 42.70 0 53.96 1 50.86 2 54.02 2
+Qwen 0.5B Winogrande 49.86 0 51.88 5 50.43 0 50.43 5 50.49 2 52.49 9
+MMLU 31.48 0 39.98 2 44.87 0 43.76 2 45.53 2 45.58 3
+Table 1: Comparison of Llama 3.1 8B and Qwen 0.5B across Winogrande, MMLU, and CommonQA under different pruning and fine-tuning regimes.
+Columns denote: (i) Baseline = original model, (ii) Pruned Only = Tale without fine-tuning, (iii) FT Only = fine-tuned without pruning, (iv) Prune $\rightarrow$ FT = prune then fine-tune, (v) FT $\rightarrow$ Prune = fine-tune then prune, (vi) (Prune $\rightarrow$ FT) $\rightarrow$ Prune = best fine-tuned-pruned model further pruned.
+Perf.
+= performance score, #D = number of deleted layers.
+Overall, these results highlight an unexpected but consistent trend: pruning with Tale does not hinder fine-tuning but instead synergizes with it.
+Pruning acts like a regularizer, simplifying the optimization landscape, and can effectively interleave with fine-tuning to create models that are both more accurate and computationally efficient.
+Pruned models fine-tune faster, require fewer parameters to adapt, and are close to or better in performance than their full counterparts.
+6 Discussion
+We summarize five key observations below from our experiments.
+1. Task dependency of layer importance
+The literature offers a discussion in the literature about layer importance.
+Some say early layers are essential (Dalvi et al., 2020); some say late Tenney et al.
+(2019); Bansal et al.
+(2023); Song et al.
+(2025).
+Our findings show that layer importance is fundamentally task specific; e.g., removing early layers reduces accuracy to near zero on commonsense reasoning tasks (Figure 12), but removing LLaMA’s layer 3 improves performance on GSM8K-hard.
+Related tasks often exhibit similar layer dependencies.
+Commonsense reasoning tasks (see Figure 12) show importance concentrated in comparable regions of the network.
+All models showed sizable accuracy boosts in mathematical reasoning tasks after from pruning minimally one to maximally three early to middle layers (e.g., LLaMA layer 3, Mistral layers 6 and 22, Lucie layer 12) (Figures 9, 10, 11).
+By contrast, knowledge-intensive tasks ARC, BoolQ, CommonsenseQA, Winogrande, and BIG-Bench) exhibit more modest improvements (though LLaMA boasted an 11% gain on BIG-Bench) and benefit from deleting later layers.
+These results may help model interpretability, as plotting performance degradation from ablating layers helps localize specific task-solving abilities in the network.
+Initial multilingual testing of Tale on Lucie, tuned for French conversational proficiency Gouvert et al.
+(2025), with bilingual versions of the same data set showed that optimal pruning was task specific rather than language specific.
+Layer redundancy results even with a model trained for a single task.
+We trained a toy transformer entirely from scratch on a single task: an in-context learning setup involving linear functions.
+Even in this controlled single-task setting, several layers proved redundant (Figure 6) or degraded performance.
+In principle Tale can combine tasks to get more general Tale optimized models.
+A LLaMA math model without layer 12 improves over baseline LLaMA on Math500 and GSM8K tasks.
+A promising method to explore is for Tale to prune models on several tasks at once with different mixtures of data to guide the pruning.
+2. Model-specific effects with Tale
+Tale affects different models differently.
+While Llama benefitted least from Tale , Lucie achieved large gains on MMLU and double-digit gains on ARC-Challenge, CommonsenseQA, BoolQ and GSM8K-hard.
+Tale conferred more modest but still substantial gains to Qwen-7B and Mistral.
+Lucie also benefitted from more substantial pruning than the other models.
+The fact that Lucie was trained on a much smaller dataset (3T tokens vs.
+15T for LLaMA and 13T for Qwen) suggests intriguing interactions between pretraining and Tale improvements.
+We suspect that models trained close to their performance ceiling (via large-scale pretraining, instruction tuning or RLHF) yield smaller gains with Tale , whereas models trained under limited objectives may benefit more.
+3. MI Layerwise Analysis.
+We use Mutual Information (MI) to investigate why selectively removing layers can improve accuracy, focusing on how information about the output evolves as it propagates through the layers.
+Unlike correlation, MI captures non-linear statistical dependencies and thus provides a more complete measure of dependence Kinney and Atwal (2014).
+We estimate MI using MINE Belghazi et al.
+(2018), which is a widely used approximation method.
+Our analysis reveals that many layers exhibit a pronounced drop in MI (Figure 8).
+Tale drops some of those layers but not all; overall Tale reduces the peaks and valleys in the graph of MI across layers.
+However, removing all layers with MI decreases yields very bad performance.
+Thus, some downward shifts in MI across adjacent layers seems necessary for the model (for more details see Appendix H).
+7 Conclusions
+Tale removes layers irrelevant to a given task $T$ to consistently yield performance above the base model on $T$ and far above the state of the art in pruning without retraining.
+Tale also reduces computation costs.
+It also profitably interacts with further training or fine tuning, further increasing task specific performance.
+Tale is a generic strategy and can prune at many levels: base pre-trained models, instruction-tuned models, fine-tuned, and post-trained models with RLHF.
+Tale can benefit high-throughput applications with time constraints (e.g.
+in multi-agent systems with task-specific agents or interactive AI assistants) and organizations facing trade-offs between model capability and computational costs to use large language models at scale.
+Limitations
+While TALE demonstrates consistent gains across a broad range of models and tasks, we note several limitations.
+First, TALE operates at the level of entire transformer layers.
+This design choice prioritizes simplicity, transparency, and compatibility with existing checkpoints, but it does not exploit finer-grained structure such as attention heads, blocks, or token-level adaptivity.
+More granular structured pruning or adaptive computation methods may provide complementary benefits, particularly when retraining or architectural modification is permissible.
+Second, TALE performs task-specific layer selection using a held-out optimization split.
+As a result, the resulting pruned architectures are specialized to individual tasks rather than universally optimal across tasks.
+While this specialization aligns with deployment scenarios where the target task is known in advance, it may be less suitable in settings that require a single model to perform well across many heterogeneous tasks without reconfiguration.
+Third, TALE relies on a greedy elimination procedure and a stopping tolerance hyperparameter to determine when further layer removal becomes detrimental.
+Although we observe stable behavior across random seeds and alternative data splits, more principled global optimization strategies or adaptive stopping criteria could further improve robustness and are left for future work.
+Finally, our empirical comparisons focus on training-free, layer-level pruning methods that preserve the original model architecture.
+Approaches that rely on retraining, block-level restructuring, or task-adaptive control flow address a different point in the design space and are therefore not directly comparable within our experimental scope.
+Overall, these limitations reflect deliberate design choices rather than deficiencies of the approach.
+We view TALE as complementary to retraining-based and fine-grained pruning methods, and believe that combining task-aware layer selection with such approaches is a promising direction for future research.
+Appendix A Implementation Details
+Hardware.
+All experiments were conducted on 1 NVIDIA A100 GPU with 80GB memory.
+Models.
+We applied Tale to five open-weights LLMs of varying scales: Qwen2.5-0.5B-Instruct, Qwen2.5-7B-Instruct, Lucie-7B-Instruct, Mistral-7B-Instruct, and Llama-3.1-8B-Instruct.
+Datasets for Tale pruning.
+The greedy layer-pruning algorithm was evaluated across nine widely used benchmarks covering reasoning, commonsense, and knowledge-intensive tasks: ARC-Challenge, ARC-Easy, MMLU, Winogrande, GSM8K, MATH500, CommonQA, BIG-Bench, and BoolQ.
+Pruning setup.
+At each iteration, Tale evaluates all candidate single-layer deletions with respect to validation accuracy.
+The pruning threshold was defined as the baseline accuracy -8% of the full model, ensuring that pruning never reduces performance relative to the original unpruned model.
+The iterative procedure terminates once no further layer removals satisfy this criterion.
+Fine-tuning setup.
+For fine-tuning experiments, we focused on Winogrande and MMLU.
+We employed LoRA with rank 64, a batch size of 4, and the optimizer paged_adamw_32bit.
+A cosine learning rate scheduler was used, and models were trained for 10 epochs.
+A.1 Data splits and leakage prevention
+For each task, we explicitly separate the data used for layer optimization from the data used for reporting results.
+Let $D$ denote the full dataset.
+We partition $D$ into two disjoint subsets: a training set $D_{\text{train}}$ and an evaluation set $D_{\text{eval}}$.
+From $D_{\text{train}}$, we further sample a small held-out subset $D_{\text{opt}}\subset D_{\text{train}}$, which is used exclusively by TALE to guide layer elimination.
+At no point does TALE access $D_{\text{eval}}$ during optimization.
+Final accuracy is computed only on $D_{\text{eval}}$.
+When official test sets are unavailable, $D_{\text{eval}}$ corresponds to a held-out validation split that is strictly disjoint from $D_{\text{opt}}$.
+Evaluation.
+The LM-Eval methodology presents a significant limitation: it selects the answer with the highest probability among the provided options rather than assessing what the model would actually generate.
+This approach ignores hallucination behavior and systematically inflates scores; for example, in a two-choice setting, a hallucinated answer still has a 50% chance of being counted as correct.
+Furthermore, LM-Eval often assigns relatively high scores to weak models, compressing performance differences and making stronger approaches appear only marginally better despite substantial real-world gains.
+This produces a misleading picture of model capability, as high LM-Eval results do not guarantee that a model will produce correct, coherent outputs in practice.
+For these reasons, we relied primarily on Decoder Eval that measures actual accuracy based on the model’s generated outputs, which we implemented for each task.
+Prompting.
+For zero-shot and few-shot evaluation, we used task-specific prompts. Below we show the prompt used for datasets, consisting of a system instruction :
+ARC-E & ARC-C System Prompt You are a Science expert assistant.
+Your task is to answer multiple-choice science questions at grade-school level.
+Each question has four answer choices, labeled A, B, C, and D.
+For each question: - Carefully read the question and all answer choices.
+- Select the single best answer from the options (A, B, C, or D).
+- Respond only with the letter of the correct answer, and nothing else—no explanation or extra words.
+Be precise and consistent: Only the answer letter.
+Bigbench System Prompt "You are a boolean expression evaluator.
+You must respond with exactly one word: either ’True’ or ’False’.
+Do not provide explanations, steps, or any other text.
+Only respond with ’True’ or ’False’."
+BOOLQ System Prompt "You are a helpful assistant that answers True/False questions based on given passages.
+Read the passage carefully and determine if the question can be answered as True or False based on the information in the passage.
+"Respond with only ’A’ for True or ’B’ for False."
+CommonQA System Prompt "You are a helpful assistant that answers multiple-choice questions requiring commonsense knowledge and reasoning.
+Read each question carefully and select the most logical answer from the given options based on common knowledge and reasoning.
+Respond with only the letter of your chosen answer (A, B, C, D, or E)."
+GSM8K System Prompt "You are a math problem solver.
+Solve the given math problem step by step.
+" "Show your complete reasoning and calculations.
+" "At the end, write your final answer after ’####’ like this: #### [your final numerical answer]""
+MMLU System Prompt "You are a helpful assistant that answers multiple-choice questions across various academic subjects including humanities, social sciences, STEM, and professional fields.
+Read each question carefully and select the best answer from the given options.
+Respond with only the letter of your chosen answer (A, B, C, or D)."
+MATH500 System prompt You are a careful math problem solver.
+Show complete step-by-step reasoning and all calculations needed to arrive at the answer.
+Use clear, numbered or labeled steps so the reasoning is easy to follow.
+IMPORTANT (formatting): After the full reasoning, write the final answer on a new line by itself in exactly this format: ####
+integer
+<integer> must be digits only, optionally with a leading “-” for negatives (e.g., -7).
+Do not add words, punctuation, units, or commentary on the same line as the #### line.
+The #### line must be the final line of the output (nothing may follow it).
+Assume all problems expect integer answers; ensure the final line contains a single integer.
+A.2 Layer Removal Implementation
+Layer pruning was implemented through a custom model wrapper that reconstructs the architecture excluding specified layers. Given delete indices $\mathcal{D}$, we create a ModifiedModel that:
+  1. 1.
+    Preserves the embedding layer, final normalization, and language modeling head from the original model
+  2. 2.
+    Constructs a new layer sequence $\mathcal{L}^{\prime}=\{l_{i}\mid i\notin\mathcal{D}\}$ by filtering out deleted layers while maintaining the order of retained layers
+  3. 3.
+    Updates the model configuration to reflect the new layer count
+The forward pass implements standard transformer computation: input embeddings are passed through the retained layers sequentially with causal attention masking, then normalized and projected to vocabulary logits.
+Position embeddings are generated automatically if not provided.
+This architecture is fully compatible with the Hugging Face training pipeline and can be directly used with LoRA fine-tuning without requiring custom training loops.
+A.3 Fine-tuning Training Details
+All fine-tuning experiments were conducted using Parameter-Efficient Fine-Tuning (PEFT) with Low-Rank Adaptation (LoRA).
+We employed 4-bit quantization using the BitsAndBytes library to reduce memory footprint during training.
+The quantization configuration used NF4 (4-bit NormalFloat) quantization type with float16 compute dtype, without nested quantization.
+LoRA Configuration:
+We applied LoRA to all linear layers in the model with the following hyperparameters: rank $r=64$, alpha $\alpha=16$, and dropout rate of $0.1$.
+These parameters were kept consistent across both full and pruned models to ensure fair comparison.
+Optimization Settings:
+Training was performed for 10 epochs using the paged AdamW optimizer (32-bit) with a learning rate of $2\times 10^{-4}$ and weight decay of $0.001$.
+We used a cosine learning rate schedule with a warmup ratio of $0.03$.
+Gradient clipping was applied with a maximum gradient norm of $0.3$.
+The effective batch size was 60 (per-device batch size of 2 with gradient accumulation steps of 30).
+Gradient checkpointing was enabled to reduce memory consumption during training.
+Data Processing:
+All sequences were truncated or padded to a maximum length of 300 tokens.
+We used right-side padding with a special padding token (ID: 128004).
+Packing was disabled to maintain sequence boundaries, and we removed unused columns from the dataset.
+The dataloader used 4 workers with the last incomplete batch dropped to ensure consistent batch sizes.
+Hardware and Implementation:
+All experiments were conducted on NVIDIA A100 GPUs.
+We used mixed-precision training without fp16 or bf16 enabled at the trainer level, relying instead on the 4-bit quantization for memory efficiency.
+Training logs were reported to TensorBoard every 25 steps.
+Dataset LLaMA 3.1 8B (zero-shot) Qwen 2.5 7B (zero-shot)
+Baseline Best Model BSBA Baseline Best Model BSBA
+Perf. Perf. $\pm$ Std #D Sp. Perf. #D Sp. Perf. Perf. $\pm$ Std #D Sp. Perf. #D Sp.
+ARC-Easy 89 90.4 $\pm$ 0.14 5 -14.6% 88.8 8 -23.5% 90.04 93.2 $\pm$ 0.16 3 -10.0% 90.08 7 -30.3%
+ARC-Challenge 79.4 80.6 $\pm$ 0.18 4 -11.7% 77.6 7 -20.5% 86.55 92.00 $\pm$ 0.15 2 -6.7% 86.55 6 -19.9%
+BoolQ 85.4 85.9 $\pm$ 0.12 3 -8.8% 85.4 7 -17.6% 81.90 83.90 $\pm$ 0.14 4 -13.3% 82.70 5 -23.2%
+MMLU 48.8 53.8 $\pm$ 0.22 1 -2.9% 50.2 9 -26.4% 68.10 71.00 $\pm$ 0.19 5 -16.6% 68.13 6 -19.9%
+CommonQA 72.9 73.8 $\pm$ 0.15 3 -8.8% 73.10 6 -17.6% 80.30 84.40 $\pm$ 0.17 2 -6.6% 80.50 6 -19.9%
+Winogrande 53.8 54.1 $\pm$ 0.20 4 -11.7% 53.83 12 -32.2% 62.04 67.25 $\pm$ 0.21 3 -10.0% 62.19 6 -19.9%
+BIG-Bench 77.2 85.6 $\pm$ 0.25 5 -14.4% 76.4 11 -32.2% 79.20 81.60 $\pm$ 0.14 6 -19.9% 81.60 6 -19.9%
+GSM8K-HARD 39.0 59.0 $\pm$ 0.30 1 -2.9% 39.4 4 -11.7% 43.80 61.80 $\pm$ 0.28 2 -43.6% 43.99 5 -17.6%
+Math500 25.4 28.2 $\pm$ 0.18 2 -6.0% 27.4 3 -9.1% 31.00 38.20 $\pm$ 0.22 2 -6.6% 32.10 4 -13.3%
+Dataset Lucie 7B (zero-shot) Mistral 7B (zero-shot)
+Baseline Best Model BSBA Baseline Best Model BSBA
+Perf. Perf. $\pm$ Std #D Sp. Perf. #D Sp. Perf. Perf. $\pm$ Std #D Sp. Perf. #D Sp.
+ARC-Easy 74.4 75.8 $\pm$ 0.15 6 -18.1% 73.8 8 -23.5% 83.8 85.6 $\pm$ 0.14 5 -15.4% 82.8 9 -27.7%
+ARC-Challenge 46.0 51.45 $\pm$ 0.22 7 -22.1% 48.8 11 -33.1% 76.2 79.1 $\pm$ 0.18 6 -18.5% 76.2 8 -24.6%
+BoolQ 53.0 74.0 $\pm$ 0.30 5 -17.2% 63.0 19 -54.2% 81.3 84.4 $\pm$ 0.16 4 -18.5% 80.8 5 -27.7%
+MMLU 13.0 54.0 $\pm$ 0.35 8 -24.1% 15 22 -60.2% 39.4 40.8 $\pm$ 0.20 2 -6.2% 39.0 8 -24.6%
+CommonQA 54.2 68.6 $\pm$ 0.27 3 -9.1% 54.6 17 -48.2% 61.0 64.4 $\pm$ 0.19 4 -12.3% 61.6 7 -21.5%
+Winogrande 51.6 53.1 $\pm$ 0.21 5 -27.1% 53.0 15 -45.2% 53.2 54.3 $\pm$ 0.16 10 -30.7% 52.4 13 -40.0%
+BIG-Bench 67.4 75.0 $\pm$ 0.25 9 -27.1% 71 15 -45.1% 70.4 75.4 $\pm$ 0.22 9 -28.0% 72.6 11 -33.8%
+GSM8K-HARD 32 39.0 $\pm$ 0.28 1 -3.1% 37 3 -9.1% 24 33 $\pm$ 0.24 2 -6.2% 26.1 4 -12.3%
+Math500 21.0 26.1 $\pm$ 0.20 2 -6.0% 25.1 3 -9.1% 19 28 $\pm$ 0.23 1 -3.1% 18.8 4 -12.3%
+Table 2: Robustness study of the proposed layer-dropping method across multiple language models under zero-shot evaluation.
+For each dataset and model, results are reported over five random seeds to account for variability in decoding and sampling.
+We present the baseline model accuracy and the accuracy of the best pruned configuration, along with their corresponding standard deviations computed across the 5 seeds.
+The table also includes the number of transformer layers removed in the best-performing configuration (#D) and the resulting inference speedup (Sp.) expressed as the percentage of total TFlops saved during evaluation.
+All experiments use 10% of the training split for optimization and evaluate on the respective test sets.
+Bold values indicate the highest mean accuracy for each dataset.
+Appendix B Robustness Study on TALE
+See Tables 2 and 17.
+Appendix C Ablation study on validation Set of Pruning
+We analyze the effect of validation set size on TALE’s layer selection.
+Table 3 reports the specific layers dropped for different validation set sizes across three tasks (ARC-Easy, MMLU, GSM8K) and two models (Llama 3.1 8B, Qwen 2.5 7B).
+Model Val Size Task Dropped Layers
+Llama 3.1 8B 200 ARC-E {19, 20, 22, 29, 32 }
+MMLU { 21 }
+GSM8K { 3 }
+500 ARC-E { 19, 20, 21, 29, 32 }
+MMLU { 21 }
+GSM8K { 3 }
+1000 ARC-E { 19, 20, 21, 29, 32 }
+MMLU { 21 }
+GSM8K { 3 }
+Qwen 2.5 7B 100 ARC-E { 22 , 27 , 28 }
+MMLU { 18 , 22 , 24 , 27 , 28 }
+GSM8K { 19 }
+500 ARC-E { 19, 22 , 28 }
+MMLU { 22 , 23 , 26 , 27 , 28 }
+GSM8K { 19 }
+1000 ARC-E { 19 , 22 , 28 }
+MMLU { 22 , 23 , 26 , 27 , 28 }
+GSM8K { 19 }
+Table 3: Layers removed by TALE for different validation-set sizes across three tasks. This reveals the stability of pruning decisions directly.
+Appendix D Number of parameters per layer for each model
+Model Params/Layer Layers
+LLaMA 3.1 8B 218,112,000 32
+Qwen 2.5 7B 233,057,792 28
+Mistral 7B 218,112,000 32
+Lucie 7B 192,946,176 32
+Qwen 2.5 0.5B 14,912,384 24
+Table 4: Model parameter counts comparison showing parameters per layer and total number of layers.
+Appendix E Practical computing savings and scaling
+We quantify TALE’s inference-cost reduction by measuring TFLOPs (tera-FLOPs) drop per removed layer.
+Across models and tasks, removing a single transformer layer yields a mean TFLOPs reduction of $3.00\%\pm 0.20\%$.
+Because TALE removes entire layers sequentially, the total TFLOPs reduction scales essentially linearly with the number of iterations (layers removed).
+In practice this means only a few iterations are required to reach common sparsity targets: e.g., three iterations remove roughly $\approx$ 9% TFLOPs, sufficient to realize $\approx$ 10% sparsity in our settings.
+Wall-clock runtime and end-to-end cost.
+In addition to FLOPs-based estimates, we report wall-clock runtime measurements under the exact decoding settings used for evaluation.
+TALE incurs a one-time optimization cost per task, after which the pruned model is used for standard inference with reduced depth.
+For all non-reasoning benchmarks (e.g., ARC, BoolQ, MMLU, CommonsenseQA), we use greedy decoding with max_new_tokens = 1, reflecting single-step answer generation.
+Under this setting, the total wall-clock time required to complete a full TALE optimization run is approximately 1 hour per dataset on a single NVIDIA A100 GPU.
+For reasoning-heavy benchmarks (GSM8K-Hard and Math500), we use max_new_tokens = 200 to allow full chain-of-thought generation.
+Due to longer decoding sequences, the average wall-clock time for TALE on these datasets is approximately 3 hours, measured over the evaluated subset.
+Importantly, this optimization cost is incurred only once per task.
+During deployment, inference is performed using the pruned model, yielding reduced per-example latency and throughput improvements proportional to the number of eliminated layers.
+These gains are amortized over all subsequent inference calls.
+Appendix F Layer Redundancy Is Not a Multi-Task Artifact
+F.1 Training details
+Figure 6: Plots showing evolution of the predictions over layers for $f(x)=x$ for a model trained on degree 1 with $D_{\cal I},D_{\cal F}\sim{\cal U}(-1,1)$.
+We train a transformer model to perform in-context learning over the class of linear functions.
+Given a prompt consisting of a small set of input–output examples $(x_{1},g(x_{1}),\dots,x_{p},g(x_{p}),x),$ the model is tasked with predicting the value $g(x)$ for the final input $x$.
+Our architecture is a 12-layer transformer with 8 attention heads, trained from scratch exclusively on the linear function family.
+The training objective follows the standard ICL formulation introduced in Garg et al.
+(2022); Akyürek et al.
+(2022):
+$\min_{\theta}$ $\mathbb{E}_{g\sim\mathcal{D}_{\mathcal{F}}}\Bigg[\mathbb{E}_{x_{1},\dots,x_{p}\sim\mathcal{D}_{\mathcal{I}}}\sum_{i=0}^{k}\ell\!\Big(y_{i+1},$ (1)
+$\qquad f^{\theta}\!\big(x_{1},g(x_{1}),\dots,x_{i+1}\big)\Big)\Bigg]$
+where $\ell$ is the squared error loss. Here, $y_{i+1}=g(x_{i+1})$ denotes the ground-truth output of the underlying linear function $g$.
+F.2 Plots
+Figure 6
+Appendix G Intution behind Tale
+(a) ARC-Challenge
+(b) MMLU
+Figure 7: Layer-wise output performance for LLaMA models: results when generating predictions from intermediate layers 1 through 32 on two different datasets.
+Appendix H Information theory: Why pruned models might perform better.
+Our results pose a puzzle: the increase in accuracy with Tale is counterintuitive: why would removing parts of a carefully trained model lead to better performance?
+One way to explore this question is mutual information.
+Alemi et al.
+(2016); Tishby and Zaslavsky (2015) use information theory (Shannon, 1948) to analyze how neural networks learn and represent data.
+Fano and Hawkins (1961) define $\text{I}(\text{X};\text{Y})$, the mutual information between two random variables $X$ and $Y$, with the equation:
+$\text{I}(\text{X};\text{Y})$ $=\text{H}(\text{Y})-\text{H}(\text{Y}\mid\text{X})$ (2)
+$=\text{H}(\text{X})-\text{H}(\text{X}\mid\text{Y})$
+$=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\log\frac{p(x,y)}{p(x)\,p(y)}$
+where $p(x,y)$ is the joint distribution of X and Y, and $p(x),p(y)$ are their marginals and where $\text{H}(\text{X})=-\sum_{x}p(x)\log p(x)$ is the Shannon (1948) entropy.
+$\text{I}(\text{X};\text{Y})$ measures how much knowing X reduces uncertainty about Y (Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017).
+To attempt to explain why accuracy increases through task pruning we also use MI.
+A major challenge of this approach is that it requires information about true distributions, which are infeasible to compute.
+As a result, researchers typically assume a Gaussian distribution Gabrié et al.
+(2019); Gao et al.
+(2015); Park et al.
+(2024) or approximate the probe using a classifier Belinkov (2022); Alain and Bengio (2016) or an MLP Belghazi et al.
+(2018).
+These approximations can yield useful insights.
+In our case, the Gaussian assumption did not fit our datasets.
+Since we evaluate on QA tasks, we used a trainable classifier to approximate the probes and estimate $I(X^{\ell},\text{Y})$ at each layer, where $X^{\ell}$ denotes the contextualized representations at layer $\ell$ and Y denotes the target answer.
+This approximates how much information the layer $\ell$ representations contain about the answer.
+The goal is then to examine whether some layers exhibit a sharp drop in information and whether those layers coincide with the ones whose removal leads to improved performance.
+Our findings, summarized in Figure 9 and Table 9, reveal two key patterns: (i) several layers in large pre-trained transformers exhibit a pronounced drop in mutual information; (ii) removing layers dictated by Tale consistently increases the mutual information at the subsequent layer across tasks.
+Together, these results suggest that certain layers act more as bottlenecks than as contributors to task-relevant representations, providing a rationale for why pruning can lead to improved accuracy.
+Figure 8: Evolution of Mutual Information about the output through layers for Llama3-8B.
+(a) ARC-Easy (Qwen 0.5B)
+(b) BoolQ (Lucie 7B)
+(c) BigBench (Llama 8B)
+Figure 9: Evolution of mutual information (MI) across transformer layers for different benchmark datasets and different models.
+Each subplot shows how information is processed and transformed as it flows through the network layers, demonstrating distinct patterns of information propagation for (a) ARC-Easy on Qwen 0.5B, (b) BoolQ on Lucie 7B, and (c) BigBench on LLaMA 8B.
+Appendix I LM-Eval as a sanity check
+LM-Eval is a commonly used multiple-choice evaluation protocol and is included here solely as a sanity check for comparability with prior work.
+The LM-Eval results reported in Table 5 are obtained by evaluating the same pruned models selected using the optimization protocol described in Appendix A.1.
+Dataset LLaMA 3.1 8B 0-shots
+Baseline Best Model BSBA
+Perf. Perf. #D Sp. Perf. #D Sp.
+BoolQ 82.4 87.63 3 - 85.62 7 -
+Hellaswag 52.5 55.5 3 - 54.5 5 -
+COMMONQA 77.2 81.61 6 - 80.27 7 -
+WINOGRANDE 75.92 78.93 4 - 76.59 5 -
+Table 5: Results of LLaMA 3.1 8B across nine benchmarks. All tested on 0-shot and evaluated with lm eval
+Appendix J A tunable metric for finding accuracy vs. speed up optimization
+To systematically select among these candidates according to user priorities, we propose the Accuracy–Efficiency Harmonic Mean (AE-HM):
+$r_{A}$ $=\frac{\text{Acc}(\text{Model})}{\text{Acc}(\text{Baseline})},$ (3)
+$\text{AE-HM}(\text{Model})$ $=\frac{(1+\lambda^{2})r_{A}\,S}{\lambda^{2}S+r_{A}}$
+$=\frac{1+\lambda^{2}}{\frac{\lambda^{2}}{r_{A}}+\frac{1}{S}}$
+where $S$ denotes the relative inference speedup and $\lambda$ controls the relative importance of accuracy versus efficiency.
+The user can set AE-HE’s parameter $\lambda$ to desired specifications: if $\lambda>1$, we prioritize $r_{A}$; if $\lambda<1$ we prioritize Speedup.
+By computing AE-HM for candidate models, we can automatically identify the model with the highest score for a given task or a set of tasks given a particular AE-HM parameter setting:
+$M_{\text{best-compromise}}=\arg\max_{i}\text{AE-HM}(M_{i})$ (4)
+Figure 10: Relative Gain comparison across datasets. LLaMA $\beta=3$
+Appendix K Results
 (a) ARC-Easy
@@
 The $\star$ denotes the best-performing layer drop configuration, while the $\square$ highlights the Best Speed up with at least Baseline Accuracy (BSBA) configuration.
-Dataset LLaMA 3.1 8B (zero-shot) Qwen 2.5 7B (zero-shot)
-Baseline Best Model BSBA Baseline Best Model BSBA
-Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
-ARC-Easy 87.00 90.55(+4.08% $\uparrow$) 5 1.3 87.82 8 1.4 90.04 94.40(+4.84% $\uparrow$) 3 1.2 90.08 7 1.5
-ARC-Challenge 75.86 78.62(+3.63% $\uparrow$) 4 1.3 76.90 7 1.4 86.55 92.00(+6.45% $\uparrow$) 2 1.3 86.55 6 1.4
-BoolQ 85.00 86.20(+1.40% $\uparrow$) 3 1.1 85.70 7 1.4 81.90 83.90(+2.44% $\uparrow$) 4 1.3 82.70 5 1.2
-MMLU 54.87 59.90(+9.17% $\uparrow$) 1 1.1 54.87 9 1.4 68.10 71.00(+4.26% $\uparrow$) 5 1.2 68.13 6 1.3
-CommonQA 72.20 75.30(+4.29% $\uparrow$) 3 1.2 73.10 6 1.3 80.30 84.40(+5.11% $\uparrow$) 2 1.1 80.50 6 1.3
-Winogrande 53.83 56.67(+5.28% $\uparrow$) 4 1.3 53.83 12 1.5 62.04 67.25(+8.40% $\uparrow$) 3 1.2 62.19 6 1.5
-BIG-Bench 75.20 83.60(+11.17% $\uparrow$) 5 1.3 75.20 11 1.6 79.20 81.60(+3.03% $\uparrow$) 6 1.5 81.60 6 1.5
-GSM8K-HARD 15.07 37.08(+146.05% $\uparrow$) 1 1.1 17.31 6 1.4 7.8 26.8(+243.58% $\uparrow$) 2 1.1 8.99 5 1.2
-Math500 20.50 26.00(+26.83% $\uparrow$) 1 1.1 26.00 3 1.2 18.00 27.00(+50.0% $\uparrow$) 2 1.1 21.00 4 1.2
-Dataset Lucie 7B (zero-shot) Mistral 7B (zero-shot)
-Baseline Best Model BSBA Baseline Best Model BSBA
-Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
-ARC-Easy 72.45 76.55(+5.66% $\uparrow$) 6 1.3 72.55 13 1.8 81.02 83.45(+4.23% $\uparrow$) 5 1.3 81.09 9 1.5
-ARC-Challenge 48.00 53.79(+12.06% $\uparrow$) 7 1.4 51.38 11 1.8 72.20 74.83(+3.64% $\uparrow$) 6 1.4 72.41 8 1.4
-BoolQ 53.70 77.50(+44.32% $\uparrow$) 5 1.3 60.60 19 2.8 80.36 83.20(+3.53% $\uparrow$) 6 1.2 80.60 10 1.5
-MMLU 21.36 42.98(+101.2% $\uparrow$) 8 1.4 22.73 22 3.3 52.73 57.81(+9.63% $\uparrow$) 2 1.0 52.91 8 1.4
-CommonQA 55.50 69.70(+25.59% $\uparrow$) 3 1.5 57.10 17 2.6 57.32 61.40(+7.12% $\uparrow$) 4 1.2 57.40 7 1.3
-Winogrande 54.20 57.80(+6.64% $\uparrow$) 5 1.2 54.30 15 1.9 52.55 58.80(+11.53% $\uparrow$) 10 1.6 53.43 13 1.8
-BIG-Bench 69.60 77.20(+9.84% $\uparrow$) 9 1.2 72.00 15 1.6 70.00 76.40(+9.14% $\uparrow$) 9 1.3 72.80 11 1.4
-GSM8K-HARD 14.20 17.80(+25.35% $\uparrow$) 1 1.1 17.40 3 1.1 11.24 19.10(+69.92% $\uparrow$) 2 1.1 15.73 4 1.2
-Math500 19.00 27.00(+42.11% $\uparrow$) 2 1.1 26.00 3 1.2 8.00 16.00(+100% $\uparrow$) 1 1.0 10.00 4 1.2
-Dataset Qwen 2.5 0.5B (zero-shot)
-Baseline Best Model BSBA
-Perf. Perf. #D Sp. Perf. #D Sp.
-ARC-Easy 40.00 60.91(+48.49% $\uparrow$) 3 1.1 48.36 5 1.4
-ARC-Challenge 35.52 40.34(+13.57% $\uparrow$) 1 1.1 37.24 4 1.5
-BoolQ 62.30 67.20(+7.87% $\uparrow$) 5 1.4 66.20 6 1.5
-MMLU 31.48 39.97(+26.96% $\uparrow$) 2 1.1 33.90 5 1.4
-CommonQA 42.40 49.10(+15.80% $\uparrow$) 2 1.3 44.00 3 1.4
-Winogrande 49.86 51.88(+4.51% $\uparrow$) 5 1.3 49.87 17 3.9
-BIG-Bench 72.40 73.60(+1.66% $\uparrow$) 2 1.2 73.60 2 1.2
-GSM8K-HARD 6.74 11.24(+66.77% $\uparrow$) 1 1.2 8.99 2 1.2
-Math500 8.00 12.00(+50% $\uparrow$) 1 1.1 9 2 1.1
-Table 1: Performance comparison across language models under 0-shot evaluation.
-We report accuracy (Perf.), number of dropped layers (#D), and relative inference speedup (Sp.).
-Percentage gain = $\frac{\text{Acc}_{\text{Best}}-\text{Acc}_{\text{Baseline}}}{\text{Acc}_{\text{Baseline}}}\times 100$.
-Best accuracy is highlighted in bold; BSBA shows balanced trade-offs.
-Iterative pruning trajectories.
-Figure 2 visualizes the iterative layer-pruning process for LLaMA 3.1 8B.
-Each curve tracks accuracy as layers are progressively removed.
-The $\star$ denotes the best-performing pruned model (Best), while the $\square$ highlights the Best Speedup with Baseline Accuracy (BSBA) model—the pruned configuration achieving maximum inference speedup without falling below baseline accuracy.
-From these trajectories, three consistent patterns emerge: (i) Tale identifies compressed models that outperform the original across diverse tasks, with $\star$ markers lying strictly above baseline.
-(ii) Accuracy improvements persist across multiple pruning steps before diminishing returns, showing that substantial redundancy exists even in carefully tuned pretrained models.
-(iii) Pruning dynamics are task-specific: datasets such as ARC-Easy and MMLU tolerate deeper pruning while continuing to improve, whereas reasoning-heavy tasks like GSM8K-Hard converge earlier, reflecting heterogeneous layer importance across domains.
-Best vs.
-BSBA models.
-Table 1 compares baseline models against their pruned counterparts under both Best and BSBA configurations.
-Across all benchmarks, the Best models yield consistent accuracy gains—up to $146\%$ (LLaMA 8B on GSM8K-Hard), $101\%$ (Lucie 7b on MMLU) and $244\%$ (Qwen 7B on GSM8k-Hard)—while also delivering moderate speedups.
-BSBA models, by construction, trade smaller gains in accuracy for more aggressive speedups, offering practical operating points when inference cost is the dominant concern.
-Few-shot setting.
-We tested Tale under the few-shot regime for Lucie and LLaMA models (Appendix Tables 45).
-Few-shot prompting improves baselines on reasoning tasks such as GSM8K and Math500, yet Tale-pruned variants still achieve higher accuracy in nearly all settings.
-This shows that pruning-induced improvements are largely complementary to gains from in-context learning.
-Takeaways.
-Tale consistently uncovers Pareto-optimal models along the accuracy–efficiency frontier.
-By balancing task fidelity with computational savings, it enables both accuracy-focused and efficiency-focused deployment.
-The observed diversity in pruning profiles across datasets underscores the importance of adaptive pruning, rather than one-size-fits-all heuristics, for effective model compression.
-A tunable selection metric for choosing among candidate trade-offs is in Appendix E.
-5.1 Tale and Fine-tuning: How Does Pruning Interact with Training?
-A natural question is whether pruning layers before or after fine-tuning harms the model’s ability to learn.
-Intuitively, one might expect that removing layers reduces representational capacity and thus limits downstream fine-tuning performance compared to baseline instruct-tuned models.
-Surprisingly, our experiments show the opposite: Tale not only preserves fine-tuning efficacy but in several cases improves both accuracy and efficiency.
-We systematically explored four settings: (i) fine-tuning the base model (FT), (ii) applying Tale after fine-tuning (FT $\rightarrow$ Tale ), (iii) pruning first and then fine-tuning (Tale $\rightarrow$ FT), and (iv) pruning first then finetuning and then again pruning (BASE $\rightarrow$ Tale $\rightarrow$ FT $\rightarrow$ Tale ).
-Across reasoning and knowledge benchmarks, we consistently observed moderate but significant gains, especially on Winogrande (Table 2).
-For example, pruning LLaMA-3.1 8B before fine-tuning reduced fine-tuning time by 2–2.5 GPU hours on an A100 (an 18.5% reduction) while simultaneously improving Winogrande performance by $2.4\%$.
-Iteratively applying pruning and fine-tuning allowed us to prune up to 8 layers achieving still higher accuracy (87.37%) than the full fine-tuned model (85.00%).
-Similarly, pruning the fully fine-tuned model yielded a 7-layer reduction while maintaining strong accuracy (86.66%).
-On knowledge-heavy tasks such as MMLU and CommonsenseQA, the gains were smaller but consistent: pruned fine-tuned models retained or slightly exceeded the accuracy of their full fine-tuned counterparts, while requiring fewer parameters to optimize.
-This suggests that pruning can act as a regularizer, simplifying the optimization landscape by removing redundant layers.
-Model Dataset Baseline Pruned Only FT Only Prune $\rightarrow$ FT FT $\rightarrow$ Prune (Prune $\rightarrow$ FT) $\rightarrow$ Prune
-Perf. #D Perf. #D Perf. #D Perf. #D Perf. #D Perf. #D
-Llama 3.1 8B Winogrande 53.83 0 56.67 4 85.00 0 87.06 4 86.74 7 87.37 8
-MMLU 54.87 0 59.90 1 63.62 0 63.49 1 64.21 2 64.01 2
-CommonQA 72.20 0 75.30 3 81.88 0 81.80 3 83.40 3 82.90 6
-GSM8K 15.07 0 37.08 3 42.70 0 53.96 1 50.86 2 54.02 2
-Qwen 0.5B Winogrande 49.86 0 51.88 5 50.43 0 50.43 5 50.49 2 52.49 9
-MMLU 31.48 0 39.98 2 44.87 0 43.76 2 45.53 2 45.58 3
-Table 2: Comparison of Llama 3.1 8B and Qwen 0.5B across Winogrande, MMLU, and CommonQA under different pruning and fine-tuning regimes.
-Columns denote: (i) Baseline = original model, (ii) Pruned Only = Tale without fine-tuning, (iii) FT Only = fine-tuned without pruning, (iv) Prune $\rightarrow$ FT = prune then fine-tune, (v) FT $\rightarrow$ Prune = fine-tune then prune, (vi) (Prune $\rightarrow$ FT) $\rightarrow$ Prune = best fine-tuned-pruned model further pruned.
-Perf.
-= performance score, #D = number of deleted layers.
-Overall, these results highlight an unexpected but consistent trend: pruning with Tale does not hinder fine-tuning but instead synergizes with it.
-Pruned models fine-tune faster, require fewer parameters to adapt, and are close to or better in performance than their full counterparts.
-Pruning finetuned models almost always improves performance.
-This shows that Tale pruning can be effectively interleaved with fine-tuning to create models that are both more accurate and computationally efficient.
-6 Information theory and why pruned models perform better.
-Our experiments show that selectively dropping layers improves accuracy for a targeted task. This seems counterintuitive: why could removing parts of a carefully trained model improve it?
-Alemi et al.
-(2016); Tishby & Zaslavsky (2015) have used information theory (Shannon, 1948) to analyze how neural networks learn and represent data.
-We can use also this tool to explain the effects of our pruning algorithm.
-Fano & Hawkins (1961) define $\text{I}(\text{X};\text{Y})$, the mutual information between two random variables $X$ and $Y$, with the equation:
-$\text{I}(\text{X};\text{Y})=\text{H}(\text{Y})-\text{H}(\text{Y}\mid\text{X})=\text{H}(\text{X})-\text{H}(\text{X}\mid\text{Y})=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\log\frac{p(x,y)}{p(x)\,p(y)}$ (1)
-where $p(x,y)$ is the joint distribution of X and Y, and $p(x),p(y)$ are their marginals and where $\text{H}(\text{X})=-\sum_{x}p(x)\log p(x)$ is the Shannon (1948) entropy.
-$\text{I}(\text{X};\text{Y})$ measures how much knowing X reduces uncertainty about Y Tishby & Zaslavsky (2015); Shwartz-Ziv & Tishby (2017).
-To apply these notions in our case, we estimate $\text{I}(\text{X}^{(l)},\text{Y})$ at each layer using trained probes as mutual information approximators.
-While this approach yields $\text{I}_{\text{probe}}(\text{X}^{(l)},\text{Y})$, an approximation of the true mutual information, probe-based estimates capture the linearly accessible task-relevant information at each layer, which is meaningful since many downstream tasks employ linear classification heads Belinkov (2022).
-This estimate together with these three observations yields an information-theoretic explanation for why architectural pruning can improve model performance.
-(i) Certain layers in large pre-trained transformers decrease mutual information between the layer’s representation and the target task (Figure 3 and Table 7).
-(ii) From Tale we select the layer $\ell$ removed at the first iteration, and compare $\mathrm{I}\!\left(X^{(\ell+1)},Y\right)$ before deletion of layer $\ell$ to its value after deletion.
-Deleting this layer always increases the mutual information at the subsequent layer on the tasks, effectively preventing the previous layer, which acts as noise, from obstructing the flow of information through the network.
-Although our findings depend on the approximative nature of probe-based MI estimation, they provide evidence that certain layers in over-parameterized transformers act as information bottlenecks, which degrade rather than refine task-relevant representations.
-By removing layers that decrease mutual information with the target task, we enable representations with higher task-relevant information to flow directly to subsequent layers using the residual connection.
-This improves the information flow through the remaining architecture, results in representations that are more predictive of the target task, and yields improved accuracy at lower computational cost.
-(a) ARC-Easy (Qwen 0.5B)
-(b) BoolQ (Lucie 7B)
-(c) BigBench (Llama 8B)
-Figure 3: Evolution of mutual information (MI) across transformer layers for different benchmark datasets and different models.
-Each subplot shows how information is processed and transformed as it flows through the network layers, demonstrating distinct patterns of information propagation for (a) ARC-Easy on Qwen 0.5B, (b) BoolQ on Lucie 7B, and (c) BigBench on LLama 8B.
-7 Discussion
-We summarize five key observations below from our experiments.
-1. Deleting later layers frequently improves performance on various tasks.
-This challenges prior claims that early layers are largely redundant and more amenable to removal, instead highlighting their essential role in preserving core task-relevant representations.
-Even deleting many late layers does not reduce accuracy below baseline, whereas removing even a single early layer is often catastrophic (see Figure 4).
-All models exhibit similar behavior.
-Figure 4: Nine benchmark tasks indicating performance after one layer is dropped from different positions in Llama3-8B.
-On the other hand, early layers appear crucial for task-solving, even though probing outputs at those layers does not yield interpretable responses.
-These results are potentially important for model interpretability.
-Plotting the performance degradation from ablating individual or consecutive layers, helps localize where specific task-solving abilities reside in the network.
-2. Task dependence of layer importance.
-Which layers improve or harm performance when removed is highly task dependent.
-Sometimes a single layer is critical: for instance, removing layer 25 of Llama-8B on CommonsenseQA causes a 50-point accuracy drop.
-Removing Llama’s layer 3 improves performance on GSM8K-hard but hurts MATH500; the revers happens when removing layer 11.
-Removing early layers (1–3) reduces accuracy to near zero on commonsense reasoning tasks (Figure 4), suggesting that certain early layers localize critical task-relevant information.
-3. Structured task-specific patterns.
-Although pruning is task-specific, related tasks often exhibit similar layer dependencies.
-Commonsense reasoning tasks (see Figure 4) show importance concentrated in comparable regions of the network.
-Mathematical reasoning tasks benefit from pruning one to three early layers (e.g., LLaMA layer 3, Mistral layers 6 and 22, Lucie layer 12), but not more (Figures 7, 8, 9).
-Commonsense and language tasks (ARC, BoolQ, CommonsenseQA, Winogrande, and BIG-Bench) benefit from deleting later layers (Tables 7, 9, 8).
-This suggests that later layers often play a decoding role for predictions into natural language, which reinforces point 1—pruning them doesn’t harm predictive capability.
-We observe stronger pruning gains in reasoning-heavy tasks under zero-shot evaluation.
-All models showed notable accuracy boosts from one or two layer deletions on mathematical reasoning (e.g., Llama’s large gain on GSM8K-hard, and double-digit gains on MATH500 for Qwen, Mistral, and Lucie).
-By contrast, knowledge-intensive tasks exhibit more modest improvements (e.g., an 11% gain for Llama on BIG-Bench).
-4. Model-specific pruning effects.
-Different models display distinct pruning behavior.
-For example, pruned Lucie achieved a 101% gain on MMLU and double-digit gains on ARC-Challenge, CommonsenseQA, BoolQ and GSM8K-hard.
-While Qwen-7B, Llama-8B and Mistral share similar architecture and scale, they had modest gains on these data sets.
-Lucie also benefitted from more substantial pruning than the other models.
-Interestingly, Lucie was trained on a much smaller dataset (3T tokens vs.
-15T for Llama and 13T for Qwen) and also tuned for French conversational proficiency Gouvert et al.
-(2025), suggesting that some layers for one of our tasks may encode multilingual-specific functionality.
-This highlights intriguing interactions between pretraining and pruning efficiency.
-One hypothesis is that models trained close to their performance ceiling (via large-scale pretraining, instruction tuning or RLHF) yield smaller pruning gains, whereas models trained under limited objectives may benefit more.
-Notably, even the large corpus trained Qwen-0.5B showed strong pruning efficiency gains.
-We experimented with producing pruned models for several tasks, and we saw that a single, general pruning improved speed up without much loss in accuracy across multiple tasks (Appendix E).
-This shows that judicious pruning can preserve generalization ability across several tasks.
-5. Tale is versatile.
-Tale can prune: base pre-trained models, instruction-tuned models (as we mainly do here), fine-tuned, and post-trained models with RLHF.
-Our results show that interleaving pruning and fine-tuning yields compounded benefits that may extend to other training regimes.
-8 Conclusions
-Tale is a generic toolkit for removing layers not needed for a certain task that yields increased performance and reduced computational costs.
-We have shown that it can profitably interact with further training or fine tuning, allowing users to further increase their models’ task specific performance.
-We have also shown that Tale can improve models both at small and larger scales.
-Tale can benefit high-throughput applications with time constraints–e.g.
-in multi-agent systems with task-specific agents, real-time recommendation engines, or interactive AI assistants.
-Tale can also help smaller companies or organizations that face critical trade-offs between model capability and computational efficiency use large language models at scale.
-Appendix A Implementation Details
-Hardware.
-All experiments were conducted on 1 NVIDIA A100 GPU with 80GB memory.
-Models.
-We applied Tale to five open-weights LLMs of varying scales: Qwen2.5-0.5B-Instruct, Qwen2.5-7B-Instruct, Lucie-7B-Instruct, Mistral-7B-Instruct, and Llama-3.1-8B-Instruct.
-Datasets for Tale pruning.
-The greedy layer-pruning algorithm was evaluated across nine widely used benchmarks covering reasoning, commonsense, and knowledge-intensive tasks: ARC-Challenge, ARC-Easy, MMLU, Winogrande, GSM8K, MATH500, CommonQA, BIG-Bench, and BoolQ.
-Pruning setup.
-At each iteration, Tale evaluates all candidate single-layer deletions with respect to validation accuracy.
-The pruning threshold was defined as the baseline accuracy of the full model, ensuring that pruning never reduces performance relative to the original unpruned model.
-The iterative procedure terminates once no further layer removals satisfy this criterion.
-Fine-tuning setup.
-For fine-tuning experiments, we focused on Winogrande and MMLU.
-We employed LoRA with rank 64, a batch size of 4, and the optimizer paged_adamw_32bit.
-A cosine learning rate scheduler was used, and models were trained for 10 epochs.
-Evaluation.
-We measured performance in terms of task accuracy on the test portions of datasets through automatic evaluation, with no human assessment involved, while inference efficiency was reported as relative speedup over the full baseline model.
-To evaluate MATH500 and GSM8K automatically without any human verification, we force the model to predict the final answer after the reasoning steps in a particular format and then calculate the accuracy.
-This could lower accuracy from what is expected; but since we used the same evaluation criteria for all the techniques and models and are interested in relative improvements or decreases in performance, the increase/decrease due to our methods from what might have been expected is justified.
-Prompting.
-For zero-shot and few-shot evaluation, we used task-specific prompts. Below we show the prompt used for datasets, consisting of a system instruction :
-ARC-E & ARC-C System Prompt You are a Science expert assistant.
-Your task is to answer multiple-choice science questions at grade-school level.
-Each question has four answer choices, labeled A, B, C, and D.
-For each question: - Carefully read the question and all answer choices.
-- Select the single best answer from the options (A, B, C, or D).
-- Respond only with the letter of the correct answer, and nothing else—no explanation or extra words.
-Be precise and consistent: Only the answer letter.
-Bigbench System Prompt ”You are a boolean expression evaluator.
-You must respond with exactly one word: either ’True’ or ’False’.
-Do not provide explanations, steps, or any other text.
-Only respond with ’True’ or ’False’.”
-BOOLQ System Prompt ”You are a helpful assistant that answers True/False questions based on given passages.
-Read the passage carefully and determine if the question can be answered as True or False based on the information in the passage.
-”Respond with only ’A’ for True or ’B’ for False.”
-CommonQA System Prompt ”You are a helpful assistant that answers multiple-choice questions requiring commonsense knowledge and reasoning.
-Read each question carefully and select the most logical answer from the given options based on common knowledge and reasoning.
-Respond with only the letter of your chosen answer (A, B, C, D, or E).”
-GSM8K System Prompt ”You are a math problem solver.
-Solve the given math problem step by step.
-” ”Show your complete reasoning and calculations.
-” ”At the end, write your final answer after ’####’ like this: #### [your final numerical answer]””
-MMLU System Prompt ”You are a helpful assistant that answers multiple-choice questions across various academic subjects including humanities, social sciences, STEM, and professional fields.
-Read each question carefully and select the best answer from the given options.
-Respond with only the letter of your chosen answer (A, B, C, or D).”
-Winogrande System Prompt You are a careful math problem solver.
-Show complete step-by-step reasoning and all calculations needed to arrive at the answer.
-Use clear, numbered or labeled steps so the reasoning is easy to follow.
-IMPORTANT (formatting): After the full reasoning, write the final answer on a new line by itself in exactly this format: ####
-integer
-<integer> must be digits only, optionally with a leading “-” for negatives (e.g., -7).
-Do not add words, punctuation, units, or commentary on the same line as the #### line.
-The #### line must be the final line of the output (nothing may follow it).
-Assume all problems expect integer answers; ensure the final line contains a single integer.
-Appendix B Number of parameters per layer for each model
-Model LLaMA 3.1 8B Qwen 2.5 7B Mistral 7B Lucie 7B Qwen 2.5 0.5B
-Parameters 218,112,000 233,057,792 218,112,000 192,946,176 14,912,384
-Table 3: Model parameter counts comparison. LLaMA 3.1 8B, Mistral 7B and Lucie 7B has 32 layers, Qwen 2.5 7B has 28 layers and Qwen 2.5 0.5B has 24 layers.
-Appendix C Intution behind Tale
-(a) ARC-Challenge
-(b) MMLU
-Figure 5: Layer-wise output performance for LLaMA models: results when generating predictions from intermediate layers 1 through 32 on three different datasets.
-Appendix D Results on few-shot learning on larger models
 Dataset Lucie 7B few-shots
@@
 Table 7: Results of LLaMA 3.1 8B across nine benchmarks. All tested on 5-shots, except gms8k and MATH500 on 8-shots
-Appendix E A tunable metric for finding accuracy vs. speed up optimization
-To systematically select among these candidates according to user priorities, we propose the Accuracy–Efficiency Harmonic Mean (AE-HM):
-$r_{A}=\frac{\text{Acc}(\text{Model})}{\text{Acc}(\text{Baseline})},\qquad\text{AE-HM}(\text{Model})=\frac{(1+\lambda^{2})r_{A}\,S}{\lambda^{2}S+r_{A}}=\frac{1+\lambda^{2}}{\frac{\lambda^{2}}{r_{A}}+\frac{1}{S}}$ (2)
-where $S$ denotes the relative inference speedup and $\lambda$ controls the relative importance of accuracy versus efficiency.
-The user can set AE-HE’s parameter $\lambda$ to desired specifications: if $\lambda>1$, we prioritize $r_{A}$; if $\lambda<1$ we prioritize Speedup.
-By computing AE-HM for candidate models, we can automatically identify the model with the highest score for a given task or a set of tasks given a particular AE-HM parameter setting:
-$M_{\text{best-compromise}}=\arg\max_{i}\text{AE-HM}(M_{i})$ (3)
-Figure 6: Relative Gain comparison across datasets. LLaMA $\beta=3$
-Appendix F Deleted Layers in each Model and Benchmark
+Appendix L Deleted Layers in each Model and Benchmark
 Dataset Best Model BSBA
@@
 Table 12: Deleted layers represented as color-ccdinline numbers. Blue = Best Model, Orange = BSBA for Mistral zero-shot.
-Appendix G Common pruned layers model
+Dataset LLaMA 3.1 8B Qwen 2.5 7B
+Baseline Best Model BSBA Baseline Best Model BSBA
+Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
+ARC-Easy 87.00 90.55 (+3.55%) 5 1.27 85.09 9 1.48 90.04 94.40 3 1.17 90.08 7 1.5
+ARC-Challenge 75.86 78.62 (+2.76%) 4 1.26 76.90 7 1.41 86.55 92.00 (+5.45%) 2 1.3 86.55 6 1.37
+BoolQ 85.00 86.20 (+1.20%) 3 1.10 83.00 8 1.36 81.9 83.9 4 1.3 80.00 7 1.45
+MMLU 54.87 59.90 (+5.03%) 1 1.05 54.87 9 1.37 68.10 71.00 5 1.24 68.13 6 1.31
+COMMONQA 72.20 75.30 (+3.10%) 3 1.21 70.10 8 1.34 80.30 84.4 2 1.12 80.50 6 1.27
+WINOGRANDE 53.83 56.67 (+2.84%) 4 1.25 53.83 12 1.47 62.04 67.25 3 1.16 62.19 6 1.48
+BIG-Bench 75.20 83.60 (+8.40%) 5 1.24 75.20 11 1.58 79.2 81.60 6 1.47 81.60 6 1.47
+GSM8K-HARD 15.07 38.02 (+22.95%) 1 1.12 17.31 6 1.35 - - - - - - -
+GSM8K* 42.15 56.56 (+14.41%) 1 48.14 3 - - - - - - -
+Table 13: Comparison of LLaMA 3.1 8B and Qwen 2.5 7B across nine benchmarks.
+We report accuracy (%), number of layers dropped, and relative inference speedup.
+Best accuracy per dataset is in bold.
+(+X%) shows accuracy improvement from baseline.
+Appendix M More on pruning and a common pruned layers model
+Figure 12: Nine benchmark tasks indicating performance after one layer is dropped from different positions in Llama3-8B.
+Appendix N General Pruning results
 Group Dataset Baseline Pruned Model speedup
@@
 Speedup is reported relative to the baseline.
+Appendix O TALE evaluation with perplexity
+Model WikiText2 LAMBADA
+Vanilla TALE Vanilla TALE
+LLaMA 3.1 8B 24.6 24.9 28.1 28.9
+Lucie 7B 46.1 36.4 52.5 43.8
+Table 15: Perplexity scores for two models across WikiText2 and LAMBADA with Vanilla and TALE (sparisty 10%) configurations.
+Model Method Sparsity WinoGr ARC-e ARC-c
+Baseline 0% 41.2 51.7 40
+LLaMA-2-7B SLEB 10% 18 (-56.3% $\downarrow$) 29 (-43.9% $\downarrow$) 28.8 (-28.0% $\downarrow$)
+TALE 10% 56 (+35.9% $\uparrow$) 62.3 (+20.5% $\uparrow$) 50 (+25.0% $\uparrow$)
+TALE 25% 51 (+23.8% $\uparrow$) 64.8 (+25.3% $\uparrow$) 47.6 (+19.0% $\uparrow$)
+Baseline 0% 42 73.0 54.9
+LLaMA-2-13B SLEB 10% 24.2 (-42.3% $\downarrow$) 43.5 (-40.4% $\downarrow$) 29.8 (-47.3% $\downarrow$)
+TALE 10% 56.4 (+34.3% $\uparrow$) 77.3 (+5.9% $\uparrow$) 64.4 (+17.1% $\uparrow$)
+TALE 25% 55.2 (+31.4% $\uparrow$) 75.3 (+3.2% $\uparrow$) 64.1 (+16.4% $\uparrow$)
+Table 16: Accuracies (%) with Decoder Eval on zero-shot tasks for LLaMA-2-7B and LLaMA-2-13B
+Appendix P Scores with relative increases in accuracy
+Dataset LLaMA 3.1 8B (zero-shot) Qwen 2.5 7B (zero-shot)
+Baseline Best Model BSBA Baseline Best Model BSBA
+Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. saved Perf. #D Sp.
+ARC-Easy 87.00 90.55(+4.08% $\uparrow$) 5 -14.6% 87.82 8 -23.5% 91.01 91.82(+0.89% $\uparrow$) 2 -10.0% 90.91 5 -30.3%
+ARC-Challenge 75.86 78.62(+3.63% $\uparrow$) 4 -11.7% 76.90 7 -20.5% 86.55 92.00(+6.45% $\uparrow$) 2 -6.7% 86.55 6 -19.9%
+BoolQ 85.00 86.20(+1.40% $\uparrow$) 3 -8.8% 85.70 7 -17.6% 84.10 86.90(+3.22% $\uparrow$) 4 -13.3% 82.70 5 -23.2%
+MMLU 54.87 59.90(+9.17% $\uparrow$) 1 -2.9% 54.87 9 -26.4% 68.10 71.00(+4.26% $\uparrow$) 5 -16.6% 68.13 6 -19.9%
+CommonQA 72.20 75.30(+4.29% $\uparrow$) 3 -8.8% 73.10 6 -17.6% 80.30 84.40(+5.11% $\uparrow$) 2 -6.6% 80.50 6 -19.9%
+Winogrande 53.83 56.67(+5.28% $\uparrow$) 4 -11.7% 53.83 12 -32.2% 62.04 67.25(+8.40% $\uparrow$) 3 -10.0% 62.19 6 -19.9%
+BIG-Bench 75.20 83.60(+11.17% $\uparrow$) 5 -14.4% 75.20 11 -32.2% 79.20 81.60(+3.03% $\uparrow$) 6 -19.9% 81.60 6 -19.9%
+GSM8K-HARD 15.07 37.08(+146.05% $\uparrow$) 1 -2.9% 35.0 4 -11.7% 7.9 27.0(+243.58% $\uparrow$) 2 -6.6% 19.1 4 -13.3%
+Math500 20.50 26.00(+26.83% $\uparrow$) 1 -2.9% 26.00 3 -8.8% 18.00 27.00(+50.0% $\uparrow$) 2 -6.6% 21.00 4 -13.3%
+Dataset Lucie 7B (zero-shot) Mistral 7B (zero-shot)
+Baseline Best Model BSBA Baseline Best Model BSBA
+Perf. Perf. #D Sp. Perf. #D Sp. Perf. Perf. #D Sp. Perf. #D Sp.
+ARC-Easy 72.45 76.55(+5.66% $\uparrow$) 6 -18.1% 72.55 13 -39.2% 81.02 83.45(+4.23% $\uparrow$) 5 -15.4% 81.09 9 -27.7%
+ARC-Challenge 48.00 53.79(+12.06% $\uparrow$) 7 -21.1% 51.38 11 -33.1% 72.20 74.83(+3.64% $\uparrow$) 6 -18.5% 72.41 8 -24.6%
+BoolQ 53.70 77.50(+44.32% $\uparrow$) 5 -17.2% 60.60 19 -54.2% 80.36 83.20(+3.53% $\uparrow$) 6 -18.5% 80.60 10 -27.7%
+MMLU 21.36 42.98(+101.2% $\uparrow$) 8 -24.1% 39.39 15 -45.2% 52.73 57.81(+9.63% $\uparrow$) 2 -6.2% 52.91 8 -24.6%
+CommonQA 55.50 69.70(+25.59% $\uparrow$) 3 -9.1% 57.10 17 -48.2% 57.32 61.40(+7.12% $\uparrow$) 4 -12.3% 57.40 7 -21.5%
+Winogrande 54.20 57.80(+6.64% $\uparrow$) 5 -27.1% 54.30 15 -45.2% 52.55 58.80(+11.53% $\uparrow$) 10 -30.7% 53.43 13 -40.0%
+BIG-Bench 69.60 77.20(+9.84% $\uparrow$) 9 -27.1% 72.00 15 -45.1% 70.00 76.40(+9.14% $\uparrow$) 9 -28.0% 72.80 11 -33.8%
+GSM8K-HARD 14.20 17.80(+25.35% $\uparrow$) 1 -3.1% 17.40 3 -9.1% 11.24 19.10(+69.92% $\uparrow$) 2 -6.2% 15.73 4 -12.3%
+Math500 19.00 27.00(+42.11% $\uparrow$) 2 -6.0% 26.00 3 -9.1% 8.00 16.00(+100% $\uparrow$) 1 -3.1% 10.00 4 -12.3%
+Dataset Qwen 2.5 0.5B (zero-shot)
+Baseline Best Model BSBA
+Perf. Perf. #D Sp. Perf. #D Sp.
+ARC-Easy 40.00 60.91(+48.49% $\uparrow$) 3 1.1 48.36 5 1.4
+ARC-Challenge 35.52 40.34(+13.57% $\uparrow$) 1 1.1 37.24 4 1.5
+BoolQ 62.30 67.20(+7.87% $\uparrow$) 5 1.4 66.20 6 1.5
+MMLU 31.48 39.97(+26.96% $\uparrow$) 2 1.1 33.90 5 1.4
+CommonQA 42.40 49.10(+15.80% $\uparrow$) 2 1.3 44.00 3 1.4
+Winogrande 49.86 51.88(+4.51% $\uparrow$) 5 1.3 49.87 17 3.9
+BIG-Bench 72.40 73.60(+1.66% $\uparrow$) 2 1.2 73.60 2 1.2
+GSM8K-HARD 6.74 11.24(+66.77% $\uparrow$) 1 1.2 8.99 2 1.2
+Math500 8.00 12.00(+50.00% $\uparrow$) 1 1.1 9.00 2 1.1
+Table 17: Performance comparison for LLaMA 3.1 8B, Qwen 2.5 7B, Lucie 7b, Mistral 7b and Qwen 2.5 0.5B under 0-shot evaluation on the 9 benchmarks with the training test split as in Table 2).
+We report accuracy (%), number of layers dropped, and relative inference speed in time.